### v1

# CrEIMBO: Cross Ensemble Interactions in Multi-view Brain Observations

Abstract
Modern recordings of neural activity provide diverse observations of neurons across brain areas, behavioral conditions, and subjects—thus presenting an exciting opportunity to reveal the fundamentals of brain-wide dynamics underlying cognitive function. Current analysis methods, however, often fail to fully harness the richness of such data as they either provide an uninterpretable representation (e.g., via “black box” deep networks) or over-simplify the model (e.g., assume stationary dynamics or analyze each session independently). Here, instead of regarding asynchronous recordings that lack alignment in neural identity or brain areas as a limitation, we exploit these diverse views of the same brain system to learn a unified model of brain dynamics. Specifically, we assume that brain observations stem from the joint activity of a set of functional neural ensembles (groups of co-active neurons) that are similar in functionality across recordings, and propose to discover the ensemble structures and their non-stationary dynamical interactions in a new model we term CrEIMBO (Cross-Ensemble Interactions in Multi-view Brain Observations). CrEIMBO identifies the composition of the per-session neural ensembles through graph-driven dictionary learning and models the ensemble dynamics as a latent sparse time-varying decomposition of global sub-circuits—thereby capturing non-stationary dynamics. CrEIMBO identifies multiple co-active sub-circuits while maintaining representation interpretability due to sharing sub-circuits across sessions. Furthermore, CrEIMBO distinguishes session-specific from global (session-invariant) computations by exploring when distinct sub-circuits are active. We demonstrate CrEIMBO’s ability to recover ground truth components in synthetic data and uncover meaningful brain dynamics, capturing cross-subject and inter- and intra-area variability, in high-density electrode recordings of humans performing a memory task.
Figure 1: Challenges and approach. Top right: Representation of the latent $K$ neural ensemble sub-circuits ($\{\bm{f}_{k}\}$) whose time-changing contributions to the full dynamics ($\bm{F}_{t}=\sum_{k=1}^{K}\bm{c}_{kt}\bm{f}_{k}$) governs the whole-brain activity. These subcircuits can either capture global task-related patterns (e.g. $\bm{f}_{1}$), capture general brain computations that vary over time and between sessions (e.g. $\bm{f}_{2}$), be active only under a specific session (e.g. $\bm{f}_{3}$), or capture both task related and general per-session computations at different period ($\bm{f}_{k}$, e.g., processing related to some visual cue presented at the beginning of the session only ). Top Left: Membership of neurons within functional neural ensembles. Bottom: Real-world data often provide distinct partial views (“sessions”) of the brain system that cannot be matched in terms of neural identity, quantity, or distribution between areas. This discrepancy may lead to biased conclusions, e.g., session 2 and session D lack representations from areas $a$ and $d$, respectively.
1 Introduction
Identifying the interactions between and within brain areas is fundamental to advancing our understanding of how the brain gives rise to behavior. Recent advances in neural recording present an exciting opportunity to study brain-wide interactions by enabling simultaneous recording of neural activity across many brain areas through multiple high-density electrodes. Such experiments, repeated over many sessions with different implantation patterns, offer multiple asynchronous recordings of the brain system, with each session encompassing hundreds of distinct neurons across regions.
Individual recording sessions provide a singular perspective of the brain dynamics, as the reinsertion of recording devices results in the capture of different subsets of neurons in each session. Thus fitting population-level models often reverts to per-session analysis (i.e., fitting each session independently), which is inherently constrained by the inability to incorporate the full set of recorded activity across all recorded brain areas across all sessions. Moreover, analyzing the data session-by-session, may hinder our ability to distinguish between computations that are session or subject specific from task related or global session invariant neural processes.
Compounding the difficulty of merging different sessions’ data are the non-linear and non-stationary nature of neural activity, as well as the presence of noise and variability across trials—both of which require additional model complexity to account for. Moreover to extract scientifically meaningful insight, the models must be interpretable in that model parameters must be readily relatable to task variables or connections between recorded units. This gap between the opportunities offered by modern neural data and the limited capabilities of current methods necessitates new approaches to leverage the richness in modern brain data and discover the fundamental neural sub-circuits governing brain activity.
One current approach used to combine multi-session data into a single model is the use of deep learning models. By training to predict all sessions’ data, deep learning models implicitly merge the datasets into the fit network weights. While these powerful models can combine information across sessions, their uninterpretable ‘black box’ nature and complexity typically hinders their ability to reveal the fundamental building blocks of neural computation. Other approaches, including tensor factorization and its variations (e.g., [3, 9]) may be limited in their ability to merge sessions of varying durations or identify the interactions between the identified neural components.
Here, we re-frame cross-session variability in neurons recorded and experimental settings not as a drawback, but as a valuable advantage. Rather than viewing the different sessions as capturing distinct neural populations, we perceive such data as providing multiple, complementary perspectives of one brain system. This approach enables a more thorough and holistic discovery of the underlying system designed to extract joint information from the entire dataset.
In particular, we hypothesize that the population dynamics at each session is driven by a common set of global sub-circuit interactions shared across sessions. To capture the non-linear and non-stationary aspects of the dynamics, we leverage recent advances in decomposed linear dynamical systems (dLDS), which enables us to model how the neural trajectories transverse a low-dimensional manifold via a time-varying combination of the sub-circuits. While we assume that the representation of the latent dynamics is shared across sessions, the measured neural ensembles are not. We thus model each session’s ensembles as independent neural representations, i.e., the membership of neurons to ensembles, that captures how the latent trajectories project into each sessions’ view of the brain.
In this work we lay out this novel model, which we term Cross Ensemble Interactions in Multi-view Brain Observations (CrEIMBO), and demonstrate its ability to capture latent dynamics in multi-view neural data. Specifically our contributions include:
  • We discover multi-regional brain dynamics through leveraging the richness of modern neural datasets while ensuring interpretability.
  • We distinguish between intra- and inter-region interactions by identifying sparse, same-area neural ensembles, deploying a specialized structured prior over per-session projections from circuit space to neural space.
  • We accurately recover ground truth components in synthetic data and discover meaningful cross-regional brain interactions underlying human brain data utilizing a shared low-dimensional circuit space across recording sessions.
2 Related Work
Over the past decade, remarkable advances in neural data acquisition technologies have enabled the measurement of numerous neurons across brain regions, subjects, and experimental settings. This has spurred follow-up work to leverage this extensive data for novel neural discoveries, including merging non-simultaneous sessions and examining cross-regional brain interactions, all while accounting for the complex, non-linear, and non-stationary nature of neural activity.
Merging non-simultaneous neural recording sessions. Existing work for combining non-simultaneous neural population recordings, here termed “sessions”, include  [26] who introduced the term “stitching” for this problem and modeled neural activity with a shared latent dynamical systems model and [24] who proposed to stitch asynchronous recordings for inferring functional connectivity by assuming shared cross-session latent input and demonstrated the advantage in observing more neurons for inferring connectivity. However, none of these methods capture multiple ($>2$) interacting neural sub-circuits active simultaneously spanning populations across multiple areas, nor do they distinguish between within- and between-area interactions, nor do they address temporal non-stationarity. In [1, 2], Bishop et al. proposed integrating the neural activity structure into inference and expanding the proposed applications of stitching to study of communication between two neural populations of brain areas and proposed a theoretical framework of when stitching is possible. In [20], the authors proposed extracting low-dimensional dynamics from multiple sessions by learning temporal co-variations across neurons and time. Both of these methods, through, assume at least a small overlap in the identity of cross-session observed neurons. LFADS [11, 25] enables stitching while addressing the need for neural overlap by finding a shared non-linear dynamical model. However, the latent dynamics may be uninterpretable in terms of interactions among neural populations, and due to its foundation on a variational auto-encoder, it may be challenging to tune.
Maintaining Interpretability in Non-Stationary Neural Dynamics Models. Real-world neural activity follows non-stationary nature due to changing environmental settings. Classical dynamical models often capture non-stationarity through various methods such as ’black box’ deep learning models (e.g. [22, 25, 29], which are powerful but are often not interpretable with respect to neural interactions; linear models, which are interpretable but not expressive enough to capture the complex nature of neural data; or switching piece-wise linear models  [15, 14, 19], which fail to enable the identification of co-active dynamics. dLDS (decomposed Linear Dynamical Systems) [16] (with its improved learning procedure in [18]), addresses this gap by generalizing the switching assumption to sparse time-changing decomposition of linear dynamical systems, thus enabling capturing co-active sub-circuits. However, while the current dLDS framework can be modified for co-processing multi-subject recordings that present variability in neuron identity and distribution, this modification lacks the structured regional prior over the loading matrix that is necessary to partition multi-area recordings such that between and within area interactions can be isolated and studied.
Identification of Functional Neural Ensembles A potential way to find interpretable loading matrix that links low to high dimensional space is through the identification of functional neural ensembles—here refereed to as groups of neurons with similar functionalities—that determine the axes in the latent space. For instance, the mDLAG model used Automatic Relevance Determination (ARD) to promote population-wise sparsity patterns of each latent variable. In [17], the authors presented an approach to address sparsity in loading matrices while accounting for multi-trial and task-conditions variability. They proposed a graph-driven regularization technique that facilitates the grouping of co-active neurons into the same ensembles and pushing apart neurons that present different functionality, while allowing for structural adjustments in ensemble configurations across different task conditions. They further enable per-trial variability in temporal activity of the ensembles. However, their model does not model the dynamic interactions between ensembles or give a closed form prior for the evolution of the ensembles’ temporal traces. While other methods, e.g., clustering approaches like  [8] or Sparse Principal Component Analysis (SPCA) [30], enable the recognition of underlying sparse groups, they either do not allow a neuron to belong to multiple groups with varying degrees of membership or do not support controlled adaptations of structural membership across sessions.
Multi-regional brain interactions. With the ability to record from multiple brain areas and the recognition of the importance of considering multi-area communication [21], recent works have delved into understanding the communication between or within brain areas. In [7, 6], the authors proposed models to identify time delays between cross-regional brain interactions, overcoming the challenge of distinguishing indirect, concurrent, and bi-directional interactions in two or more populations. However, their model focuses on analyzing sessions individually and is not intended to uncover the full underlying set of transition matrices, but rather to recognize meaningful time delays that can imply interactions. Other models addressed multi-regional interactions through a communication subspace [23] using dimensionality reduction, or by Generalized Linear Models (GLMs), with either Poisson [28] or Gaussian [28] statistics to identify functional coupling between areas. These approaches, though, are not tailored to capture non-simultaneous cross-sessions variability in neuron identity and subject processing. Other recent methods, including [10, 5, 13] discussed the potential usage of their model to multi-regional interactions, have proposed studying multi-regional interactions using switching dynamical systems approaches, where the model switches between different states over time to capture distinct periods. However, this approach cannot disentangle interactions among multiple sub-circuits active simultaneously, nor can it identify interactions at multiple temporal resolutions that occur concurrently and distinguish between these or multiple co-processed processes whose encoding occurs gradually rather than through sharp transitions.
3 Problem Formulation and Approach
Problem Formulation
Let $\{\bm{Y}^{d}\}_{d=1}^{D}$ be a set of estimated firing rates over $D$ asynchronous sessions, indexed by ${d=1\dots D}$. The neural recordings of each $d$ session, $\bm{Y}^{d}\in\mathbb{R}^{N^{d}\times T^{d}}$ captures $T^{d}$ time observations of the activity of $N^{d}$ neurons from up to $J$ distinct brain areas, thus offering a partial view of the brain system. Note that sessions may refer to the same or different subjects, can vary in duration or number of neurons, i.e. $N^{d}$ and $T^{d}$ need not be equal to $N^{d^{\prime}}$ and $T^{d^{\prime}}$ for different sessions $d,d^{\prime}$, and may differ in the sub-set of brain areas (out of $J$) they capture. Hence, these cross-session data matrices cannot be aligned into a single data array, e.g., via a tensor—hindering the direct application of existing analyses as tensor factorization.
We assume that these brain observations reflect the hidden activity $\bm{X}^{d}\in\mathbb{R}^{p\times T^{d}}$ of $p<<N^{d}$ functional neural groups that evolve and interact over time. Each of these groups encompasses a sparse set of same-area neurons, while each neuron can belong to more than one group with varying degrees of cross-group membership, encoded in the values of $\bm{X}^{d}$. The low-dimensional latent space in which the groups interact is shared across sessions and projected to each session’s observed neurons via an unknown per-session projection $g^{d}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{N^{d}}$ that captures the groups’ compositions in the visible neurons, such that $\bm{Y}^{d}=g^{d}(\bm{X}^{d})$.
The interactions between these functional groups follow a per-session non-linear and non-statioary dynamics $\bm{F}_{t}^{d}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$, where $\bm{X}^{d}_{t}=\bm{F}^{d}_{t}(\bm{x}^{d}_{t-1})$. As in general non-stationary systems cannot be fit without constraints, we follow the dLDS model and assume that these dynamics can be described by a set $\{\bm{f}_{k}\}_{k=1}^{K}$ of $K$ global time- and session-invariant “sub-circuits”. Here each sub-circuit ${\bm{f}_{k}\in\mathbb{R}^{p\times p}}$ represents a basis interaction between these groups that is reused by the system at different time-points throughout its trajectory. Each of these latent sub-circuits may capture either “global” session-invariant brain interactions and/or session- and subject-specific interactions. Moreover, these sub-circuits may be active simultaneously or intermittently, yielding a model that can flexibly fit neural trajectories that differ between sessions with a single underlying mechanism.
The problem we address in this model is identifying the set of unknown multi-regional sub-circuits, including their neuronal composition and non-stationary activation levels, from asynchronously collected observations.
CrEIMBO (Cross Ensemble Interactions in Multi-view Brain Observations)
A naive approach to identifying the unknown sub-circuits involves first projecting each session’s data into a low-dimensional space and subsequently fitting the dynamics model in the low-dimensional space. However, this two-step approach can be sub-optimal as the dimensionality reduction projections may not prioritize the same dimensions needed to preserve dynamic fidelity. An additional advantage of simultaneously identifying the ensembles and dynamics is that in the cross-session model it automatically aligns the cross-session components based on the shared dynamics prior. Such alignment is unlikely given the sensitivity of the subspace ordering in dimensionality reduction methods to each session’s distinct statistics.
We therefore present CrEIMBO, an algorithm that simultaneously fits the global sub-circuits that underlie the inter-ensemble interactions across multiple sessions. Our model is predicated on three key assumptions that 1) ensembles of neurons, rather than individual neurons, form the basic units that interact in brain dynamics, 2) it is feasible to identify functionally analogous neuronal ensembles across sessions, and 3) the full repertoire of interactions between these ensembles are governed by a global set of dynamical primitives in the form of a basis of linear dynamical systems.
Let $\bm{A}^{d}$ be a sparse matrix where each column, $\bm{A}^{d}[:,j]$, encodes the neuronal composition of the $j$-th ensemble in the $d$-th session, such that non-zero values within these columns represent the membership of neurons in the ensemble and their magnitudes reflecting the degree of membership. Since we aim to capture multi-regional interactions, we design the ensemble matrix $\bm{a}:=\bm{A}^{d}$ as a block diagonal matrix where the $j^{th}$ block, $\bm{a}^{j}\in\mathbb{R}^{n_{j}\times p_{j}}$, contains the neural ensembles for the $j$-th area. We denote $p_{j}$ as the number of ensembles within that area (note that $p=\sum_{j=1}^{J}p_{j}$) and $n_{j}:=N_{j}^{d}$ as the number of neurons in that area (i.e., the total number of neurons observed in session $d$ is $N^{d}=\sum_{j=1}^{J}n_{j}$).
Furthermore, let $\tau:=\mathrm{T}_{m}^{d}$ indicate the time duration (i.e., the number of time-points) of the $m^{th}$ trial, for $m=1\dots M_{d}$, in session $d$. In each trial, the temporal activity for the $p$ ensembles are denoted as $\bm{x}:=\bm{X}_{m}^{d}\in\mathbb{R}^{p\times\tau}$, leading to a model where the observations $\bm{y}:=\bm{Y}_{m}^{d}$ in trial $m$ are assumed to arise from the joint activity of all these ensembles, up do a normally-distributed error $\epsilon\sim\mathcal{N}(0,\sigma)$, i.e., $\bm{y}=\bm{a}\bm{x}+\epsilon.$ The evolution of the ensembles’ activities ($\bm{x}$) reflect the latent interactions between ensembles, which we model with a non-stationaly linear dynamical system $\bm{x}_{t}=\bm{F}_{t}\bm{x}_{t-1}$ for $t=1\ldots\tau$. The time-changing transition matrix $\bm{F}_{t}\in\mathbb{R}^{p\times p}$ captures both non-linear and non-stationary brain activity. A key desired property of our model is to capture the overlapping activity of multiple sub-circuits of interacting ensembles. We thus follow the dLDS concept and model the interactions ($\bm{F}_{t}$) through a time-varying sparse decomposition ${\bm{F}_{t}=\sum_{k=1}^{K}\bm{c}_{kt}\bm{f}_{k}}$, where $\{\bm{f}_{k}\}_{k=1}^{K}$ represents the $K$ global ensemble interactions, and $\bm{c}_{kt}$ are the time-varying coefficients that represent the modulation of each interaction pattern over time.
Fitting this model requires identifying the global operators $\{\bm{f}_{k}\}_{k=1}^{K}$ and the functional ensembles $\{\bm{A}^{d}\}_{d=1}^{D}$. For this, we employ an alternating approach, where we iterate between updating the global operators, $\{\bm{f}_{k}\}_{k=1}^{K}$, and infer, for each session $d$, 1) the ensemble compositions $\bm{a}:=\bm{A}^{d}$, 2) their dynamics $\bm{x}$, and 3) the sub-circuits’ temporal coefficients $\bm{c}_{t}$. These steps are iterated until the parameters converge.
Ensemble update: We update $\bm{a}$ per-row (neuron) $n$ ($\bm{a}_{n}$). Following the work of [17], we infer the sparse ensemble structures with graph-driven re-weighted $\ell_{1}$ regularization that groups together neurons with shared activity patterns, while pushing apart neurons who do not behave similarly. The graph for each area $j$, $\bm{h}^{j}\in\mathbb{R}^{n_{j}\times n_{j}}$, is calculated based on a data-driven Gaussian kernel that measures the temporal similarity between neurons, such that $\bm{h}^{j}_{n1,n2}=exp(\frac{\|\bm{y}^{j}_{n1}-\bm{y}^{j}_{n2}\|_{2}^{2}}{\sigma_{h}})$ where $\sigma_{h}$ is a hyperparameter. Specifically, for each row $\bm{a}_{n,:}$ and considering only the ensembles in region $j_{n}$ that neuron $n$ is part of,
$\widehat{\bm{a}}_{n,:}=\arg\min_{\bm{a}_{n:}}\|\bm{y}_{n:}-\bm{a}_{n:}\bm{x}_{j_{n}:}\|_{2}^{2}+\sum_{j=1}^{p}\bm{\lambda}^{d}_{n,j}|\bm{a}_{n,j}|\quad\textrm{ where }\quad\bm{\lambda}_{n,j}=\frac{\beta_{1}}{\beta_{2}+|\widehat{\bm{a}}_{n,j}|+\beta_{3}|\bm{h}_{n:}^{j}\widehat{\bm{a}}_{:j}|}.$
Here, $\bm{x}_{j_{n},:}$ is the activity of the ensembles of the $j_{n}^{th}$ region and $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are scalars that control the effect of the graph on the regularization. The graph adjacency $\bm{h}_{n:}\widehat{\bm{a}}_{:j}$ weights the sparse membership in the $j$-th ensemble based on temporal similarity. Specifically, if $\bm{h}_{n:}\widehat{\bm{a}}_{:j}$ is high (i.e., the neighbors neuron $n$ on the graph are in the current estimate of the $j$-th ensemble), $\bm{\lambda}_{n,j}$ is small, resulting in lower sparsity regularization on neuron $n$ for that ensemble, promoting it to be included. Alternatively, if this correlation is small, the regularization weight $\bm{\lambda}_{n,j}$ is large, and neuron $n$ will be more likely not to be included in ensemble $j$.
Latent state and dynamics coefficient update: The ensemble activities $\bm{x}$ and coefficients $\bm{c}$ are updated iterative for each time-point $t$ via the LASSO optimization
${\widehat{\bm{x}}_{t},\widehat{\bm{c}}_{t}=\arg\min_{\bm{x}_{t},\bm{c}_{t}}\|\bm{y}_{t}-\bm{a}\bm{x}_{t}\|_{2}^{2}+\left\|\bm{x}_{t}-\left(\sum_{k=1}^{K}\bm{f}_{k}\bm{c}_{k,(t-1)}\right)\bm{x}_{t-1}\right\|_{2}^{2}+\lambda_{x}\|\bm{x}_{t}\|_{1}+\lambda_{c}\|\bm{c}_{t}\|_{1}}$ (1)
where $\lambda_{c}\in\mathbb{R}$ and $\lambda_{x}\in\mathbb{R}$ are regularization weights.
Dynamical system update: The global ensembles interactions, $\{\bm{f}_{k}\}_{k=1}^{K}$ are assumed to be sparse (i.e. ensembles interactions are not all to all) and are identified directly by:
$\bm{F}^{\textrm{all}}_{n}=\arg\min_{{\bm{F}^{\textrm{all}}}_{j}}\|\bm{y}_{n}-\bm{F}_{n}^{\textrm{all}}\bm{cx}\|+\lambda_{F}\|\bm{F}^{\textrm{all}}_{j}\|_{1}+\sum_{{k_{1}},{k_{2}},({k_{1}}\neq{k_{2}})}\rho({\bm{f}}_{k_{1}},{\bm{f}}_{k_{2}})$ (2)
where $\bm{F}^{\textrm{all}}$ is an horizontal concatenation of all $\bm{f}$s, and $\sum_{{k_{1}},{k_{2}}}\rho(\bm{f}_{k_{1}},\bm{f}_{k_{2}})$ is a de-correlation term used to ensure that distinct $\bm{f}$s are not too similar, Above, $\bm{cx}_{t}:=(\bm{c}^{batch}_{t}\otimes[1]_{1\times p})\circ([1]_{1\times K}\otimes\bm{x}^{batch}_{t})$, and $\lambda_{F}$ is the weight on the sparsity-regularization on $\bm{F}$ and “batch” means that this operation is in practice taken on a random subset of sessions for computational complexity considerations.
Figure 2: A-E: Testing CrEIMBO over 312 iterations of random initializations and parameters (see Sec. D) reveals high correlations between ground truth and fit components for all unknowns, $\bm{Y}$, $\bm{A}$, $\bm{x}$, $\bm{F}$, $\bm{c}$ (subplots A-E respectively). F-G: A comparison between the ensemble trajectory ($\bm{X}$) and the ensemble compositions ($\bm{A}$) as identified by CrEIMBO and other approaches compared to the ground-truth. Details about the other approaches are found in Section F. Additional sessions in Figure 5. H-J: $-\log(MSE)$ between the ground truth latent dynamics, ensemble composition, and sub-circuit, to the corresponding components as identified by CrEIMBO vs. other approaches. The sub-circuits were re-ordered for each method to maximize correlation with the ground-truth circuits using Scipy’s linear_sum_assignment [4].
Figure 3: CrEIMBO identifies cross-regional neural sub-circuits underlying multi-session human brain recordings. A: Two exemplary observations ($\bm{Y}$, right) compared to their reconstruction by crEIMBO ($\widehat{\bm{Y}}$, left). B: Reconstruction performance in terms of $R^{2}$ and relative error across all sessions. C: The sub-circuit coefficients for 3 exemplary subjects. D: The standard deviation (std) of the coefficients over time across the two exemplary sessions. E: Examplary identified sparse ensemble matrices ($\{\bm{A}^{d}\}_{d=1}^{D}$) for the observations from A. F: Relative std of the coefficients across four subjects reveals usage of similar sub-circuits. G: The identified ensembles’ activity ($\bm{X}^{d}$) for the two exemplary sessions. H: The identified sub-circuits. Width indicates effect magnitude. Each circle corresponds to an ensemble with its color indicating the ensemble’s area. Black arrows near $\bm{f}_{8}$ and $\bm{f}_{11}$ highlight within-region interactions. I: Pairwise correlations between within-session sub-circuit coefficients (left, middle) and average within-session correlations across all sessions with std marked by the size of a “+” marker. J: Average and std of connection values in $f$s. K: Number of non-zero connections per sub-circuit $\bm{f}_{k}$.
4 Experiments
4.1 CrEIMBO recovers Ground Truth Components in Synthetic Data
We first assess CrEIMBO’s ability to recover ground-truth components from synthetic data. We generated $K=3$ sub-circuits represented by a set of 3 rotational matrices (Fig. 4B). These sub-circuits capture interactions among $p=3$ ensembles from $J=3$ regions ($p_{j}=1$ for all $j=1,\dots,J$). Each region consists of a random number of neurons drawn from a uniform distribution between $4$ to $9$. We finally generated the data by simulating 5 sessions with different neuron identities and varying numbers of neurons per region(Fig. 4A and D).
CrEIMBO was able to recover the ground truth components with high accuracy, as measured by correlation with the ground truth, across $312$ random noisy initialization repeats with random parameters from a range described in 1. Accurate recover included all parameters: the reconstructed observations $\bm{y}$ (Fig. 2A), ensemble composition $\bm{A}$ (Fig. 2B) and activity $\bm{x}$ (Fig. 2C), sub-circuits $\bm{F}$ (Fig. 2D), and coefficients $\bm{c}$ (Fig. 2E). Comparing CrEIMBO’s results to components identified by other methods (baselines detailed in F), we found that approaches that either separate the ensembles and dynamics identifications steps, analyze each session separately, or do not consider the localized structure of the ensembles, fail to recover the ground-truth components, i.e., the resulting fit model parameters are less correlated with the ground-truth observations (Fig. 2F-J, 5).
Next, we tested CrEIMBO on a richer synthetic example with $D=15$ sessions. Each session had a varying number of neurons (a random integer between $14$ and $19$ per region from a uniform distribution) and distinct sub-circuit coefficients (see Fig. 8 for the generated ground truth components and 6 for regional distributions and masking). The sub-circuits were set as rotational matrices, similar to the previous experiment. Over $204$ random initializations with different parameters (see Sec. D) we again found that CrEIMBO robustly identifies the ground truth latent dynamics, ensemble compositions, and sub-circuits. Moreover, CrEIMBO accurately reconstructed the observations with high correlation to the ground truth across all initializations (Fig. 8E) with low relative error (Fig. 8F).
We further tested CrEIMBO on more advanced data that included $D=40$ sessions with a maximum of $4$ regions and $p_{j}=3$ ensembles per region (Fig. 9). We found that CrEIMBO manages to recover the components with high correlation with the ground-truth (Fig 10).
4.2 CrEIMBO Discovers Multi-Regional Dynamics in Human Electrophysiology Data
We then tested CrEIMBO on human neural recordings from a high-density electrode array provided by [12]. The data consists of neural activity from $p=10$ brain areas in the human medial temporal and medial frontal lobes, recorded while subjects performed a screening task (detailed in [12]). The data encompasses $21$ subjects across $D=41$ non-simultaneous sessions. For each subject (except Subject 19), the data offers two recording sessions: 1) a “Screening” session to identify natural images that elicit strong neural responses for that subject and 2) a “Sternberg test” session to assess working memory using the images identified in the screening session. For Subject 19, the data includes only the Sternberg session. The data encompasses multiple regions with limited cross-session regional overlap, a few missing samples per session, and variability in the distribution and count of neurons over areas, with some areas having no neurons in some trials or subjects (Fig. 11, 12).
We first converted the spike-sorted units to estimated firing rate matrices by convolving the spike trains with a 30 ms Gaussian kernel (Fig. 7). Applying CrEIMBO to the firing rate data with a maximum of $K=12$ sub-circuits and $p_{j}=7$ ensembles (Fig. 13) per region, we found multi-regional motifs with variability across sessions.
We then tested CrEIMBO on the first $18$ seconds from whole sessions together with a maximum of $K=12$ sparse sub-circuits and as most $p_{j}=7$ ensembles per region (full parameter list available in Tab. 2). First we note that CrEIMBO can accurately reconstruct the data with high accuracy across all sessions (Fig. 3A, B). Interestingly, the ensemble compositions identified by CrEIMBO (Fig. 3E) presented sparse patterns with one dense ensemble for most regions that we hypothesize captures the ‘general’ mean field activity of that region (the right-most column of each diagonal block in Fig. 3E). The sparse ensembles we hypothesize capture the more specialized, nuanced functionality. The ensemble dynamics ($\bm{X}^{d}$, Fig. 3G) reveal that some ensembles are constantly active, implying that they might play an important role in the observed neural activity.
The identified set of global sub-circuits ($\{\bm{f}_{k}\}_{k=1}^{K}$) exhibits distinct localized motifs (Fig. 3H), i.e., most sub-circuits present clear trends of either sourcing or targeting the same area. This is in contrast to other approaches (Sec. F), that yield overlapping sub-circuits (Fig. 19) that emphasize only limited circuitry (Fig. 20). Interestingly, sub-circuits 10 and 12 ended up almost empty (mainly zero-ish values, Fig. 3H, 16). This highlights CrEIMBO’s ability to automatically nullify extra or redundant sub-circuits—i.e., a form of model selection—via the sparsity regularization. The number of active connections and average connection strength (Fig. 7J, K) present modest variability between sub-circuits. This underscores the model’s ability to identify distinct sub-circuits with different structural behaviors, capturing various interaction resolutions. Specifically, it can find sub-circuits of varying strengths and is not limited to only k-sparse sub-circuits. From a neuroscience perspective, this suggests that underlying sub-circuits can vary in density and the number of participating areas and ensembles.
Interestingly, most identified interactions occur between distinct areas (i.e., inter-region interactions) with either multiple ensembles from the same area affecting together ensembles from a different area (e.g., $\bm{f}_{3}$), or multiple ensembles from diverse areas converging to the same target area (e.g., $\bm{f}_{5}$). However, these source or target groups of ensembles vary across sub-circuits, hence the full repertoire of activity cannot be captured by a single. In addition to these between-area interactions, we also observed, although to a lesser degree, within-area interactions (e.g., as in $\bm{f}_{8}$ and $\bm{f}_{11}$, as marked by black arrows). These dynamics emphasize that within-area amplification or regulation dynamics do exist in addition to global cross-region computations, however seem to be less salient than cross-area communication.
When exploring the activation patterns of these sub-circuits (Fig. 3D, Fig. 17), we reveal variability between subjects and sessions. However, the usage of overall similar sub-circuits, with some sub-circuits not being used at all in any of the exemplary sessions (Fig. 3C, F, e.g., $\bm{f}_{12}$ not being used), while others are constantly used with high variability over time (e.g., $\bm{f}_{8}$ light-green). Interestingly, when exploring the within-session pairwise correlations between sub-circuits coefficients (Fig. 7I), it seems that most coefficients present very low ($<0.1$) correlations, indicating that the activity of the corresponding circuits is not similar and they might capture different cognitive processes, while other pairs of sub-circuits may show similarities in activity. This might indicate that for a specific subject or session, they may be related. For instance, a specific image may evoke some emotional response in one subject but not in another (e.g., $\bm{c}_{9}$ $\bm{c}_{11}$ in the second session of subject 10, however not in the first session of that subject nor in the average correlations across all sessions).
5 Discussion.
In this paper we introduced CrEIMBO (Cross Ensemble Interactions in Multi-view Brain Observations), a novel approach for uncovering multi-regional dynamics in neural data collected across multiple sessions. CrEIMBO addresses the challenge of integrating non-simultaneous neural recordings by leveraging joint inference and sparsity regularization to capture the underlying neural sub-circuits governing brain activity.
CrEIMBO offers several advantages over other models of whole-brain multi-session data. Chiefly, it identifies a cross-session shared latent space where non-stationary ensemble interactions are governed by a set of global sub-circuits. By structuring the time-varying dynamics in terms of a basis set of sub-circuits, CrEIMBO allows for the exploration of active sub-circuits across sessions and aiding in the identification of subject-to-subject and cross-session variability in neural activity. Additionally, constraints on the loading matrices $\bm{A}$ enable the identification of area-specific ensembles, facilitating the understanding of within-area compared to between-area interactions. Moreover, the simultaneous fitting of the ensembles and dynamics automatically aligns cross-session ensembles in terms of functionality, which is critical for studying cross-session variability. We demonstrated the efficacy of CrEIMBO through both synthetic experiments and neural recordings. On synthetic data, CrEIMBO consistently recovered ground truth components with high accuracy. In our neural recordings analysis, CrEIMBO found cross-regional motifs that span interactions across sessions and subjects.
Limitations and future work: While CrEIMBO leverages information from multiple brain sessions, the most modern Electrophysiology tools are still far from observing every neuron or region simultaneously. Consequently, and as with all dynamical systems models, obscured interactions with missing (i.e., unrecorded) neurons within the brain may limit the precision of the identified interactions. Additionally, CrEIMBO’s current implementation does not include an offset term. Adding this would increase complexity and require thorough exploration; hence, we will address it in future work. We limit CrEIMBO here to linear projections from the ensemble-space to per-session observation space. Extending this aspect of the model to more flexible, non-linear projections while maintaining interpretability is challenging and a future goal.
CrEIMBO’s effectiveness in aligning ensembles across sessions relies on the assumption that these sessions share commonalities in their active sub-circuits. If sessions operate with entirely different sub-circuits, cross-session ensemble alignment is not guaranteed. Another limitation is that we assume Gaussian observation statistics, i.e., spiking data must be converted to rates to use CrEIMBO. Future research could incorporate non-Gaussian likelihoods, such as Poisson for spike counts. Finally, CrEIMBO rests on the mathematical foundation of dictionary learning, which can be computationally intensive at higher dimensions. To enhance computational efficiency, we plan to implement parallel processing, data batching, and GPU processing in future iterations.
We also look to leverage CrEIMBO to identify pathological or cognitive states during which specific sub-circuits manifest exclusively (e.g., neural circuitry changes during epileptic attacks). Another extension for the current method includes extending the global cross-session circuits set to enable small, controlled structural adaptations of the sub-circuits between sessions, e.g., through an $\ell_{2}$ regularization on the distance between cross-session corresponding sub-circuits $\|\bm{f}_{k}^{d}-\bm{f}_{k}^{d^{\prime}}\|_{F}^{2}$.
Acknowledgments
N.M. was supported by the Kavli Discovery Award of the Kavli Foundation.
Appendix A Notations
Symbol Description
$D$ Number of sessions (encompassing both same and different subjects)
$M_{d}$ Number of trials for each session $d$
$\bm{Y}_{m}^{d}$ Neural recordings firing rate estimation for trial $m$ and session $d$
$N_{m}^{d}$ Number of neurons observed in trial $m$ of session $d$
$\mathrm{T}_{m}^{d}$ Number of time points observed in trial $m$ of session $d$
$\{\bm{Y}_{m}^{d}\}_{m,d}$ Overall set of observations
$\bm{A}_{m}^{d}$ The neuronal composition of each sparse ensemble in trial $m$ and session $d$
$\bm{A}_{\text{block}}$ Block of $\bm{A}$ representing neural dynamics in a specific brain area
$n_{j}$ Total neurons in area $j$
$p_{j}$ Maximum number of ensembles discernible within area $j$
$p$ Total number of ensembles across all areas
$J$ Total number of distinct brain areas
$n$ Total neurons involved across all trials or sessions
$\bm{x}_{t}$ State vector at time $t$
$\bm{F}_{t}$ Transition matrix at time $t$, representing ensemble interactions
$K$ Number of global interacting units
$\bm{c}_{kt}$ Influence of the $k$-th global interaction at time $t$
$\bm{f}_{k}$ Basic linear system representing the $k$-th global interaction
$\tau=\mathrm{T}_{m}^{d}$ Duration of the trial or session
Appendix B Differences from dLDS
While our proposed model draws inspiration from dLDS [16], it significantly diverges in its approach, particularly in ensemble identification and the details thereof. Below we list the key differences:
  1. 1.
    Incorporation of sparsity constraints on the operators $f$.
  2. 2.
    Extension to multi-sessions through shared $f$s and distinct projections ($D$) instead of coefficients ($c$).
  3. 3.
    Imposition of regional structure on $D$.
  4. 4.
    Integration of graph-driven sparsity on $D$.
  5. 5.
    Modifications to the inference procedure inspired by linocs.
  6. 6.
    Ability to capture inter-subject variability.
  7. 7.
    Adaptation to handle variations in trial duration, fluctuations in neuron count, and changes in distributions between sessions.
Appendix C Differences from SiBBlInGS
Although our proposed model draws inspiration from SiBBlInGS [17] for ensemble identification, it introduces several fundamental differences:
  1. 1.
    SiBBlInGS does not consider interactions between ensembles, whereas our model incorporates ensemble interactions.
  2. 2.
    SiBBlInGS does not address differences in the number of neurons between subjects, whereas our model accounts for this variability.
  3. 3.
    SiBBlInGS lacks consideration for region-specific structure, focusing primarily on a single block of $D$, whereas our model imposes region-specific structure on $D$.
  4. 4.
    SiBBlInGS does not account for the evolution of the ensemble activity space, whereas our model incorporates this aspect.
Appendix D Parameters Range for Synthetic Data Experiment
The full set of parameters for the synthetic experiments are available in Table 1.
Table 1: Parameter values for synthetic experiments
Parameter Value (range) Explanation
same_c False Indicates if to use shared sub-circuits coefficients across sessions.
step_f 0.1 - 0.5 Gradient descent step size for updating $\bm{f}$
GD_decay 0.99 - 0.999 Decay over iterations of gradient-descent step size.
max_error 1e-09 The error threshold to stop training the model.
max_iter 500 Threshold on the maximum number of iterations.
include_D True Indicates the inclusion of a projection to a latent space.
step_D $10^{-5}-10^{-2}$ Range of Gradient Descent Step size for $\bm{A}$.
seed 0 - 4000 Random seed to choose from
normalize_eig True , False Indicates if to normalize the sub-circuits by
start_sparse_c False , True Whether to initialize $c$ to be sparse.
sparse_f True Whether to apply sparsity on the sub-circuits
num_gradient_steps 1 -3 Number of gradient steps in an iteration.
add_avg True , False Whether to add a moving average to an iteration.
sparsity_on_f_max 40-60 Percentile of sparsity applied on each $\bm{f}$
take_multiple_gd False, True Indicates if to take multiple steps of gradient descent.
D_graph_driven True Indicates that D is inferred with graph-driven way.
infer_x_c_together False , True Indicates if the inference of $\bm{x}$ and $\bm{c}$ occurred simultaneously.
include_mask True Indicates the inclusion of a mask on $\bm{A}$
norm_D_cols True Indicates the normalization of columns of $\bm{A}$
lambda_D 0.1 - 0.3 Regularization weight on updating $\bm{A}$
step_D_decay 0.99 - 0.9999 Decay rate for updating $\bm{A}$ parameters
num_regions 3 Number of brain regions.
lambda_x 0 - 0.1 Sparsity term on the latent dynamics.
latent_dim 3 Latent dimension.
noise_level 0.2-0.6 Standard deviation of noise added to the data.
num_subdyns 3 Number of sub-circuits.
lasso_solver ’spgl1’ Pylops solver used for $\ell_{1}$ regularization.
Appendix E Parameters for human multi-regional experiment
The full set of parameters for the human data experiment are available in Table 2.
Table 2: Parameter values for real-world human experiment
Parameter Value Explanation
same_c False Indicates if sub-circuit coefficients are shared across sessions.
step_f 0.2 Step size for gradient descent in updating $\bm{f}$.
GD_decay 0.992 Gradient descent step-size decay rate over iterations.
max_error $1\times 10^{-9}$ Maximum allowable error.
max_iter 100 Maximum number of iterations.
seed 0 Random seed.
normalize_eig True Normalize eigenvalues.
start_sparse_c False Start with sparse $\bm{c}$.
include_D True Whether to include $\bm{A}$.
step_D 0.0001 Step size for updating $\bm{A}$.
num_gradient_steps 4 Number of gradient steps.
add_avg True Add average.
sparsity_on_f_max 45 Maximum sparsity on $\bm{f}$.
take_multiple_gd False Take multiple gradient descents.
D_graph_driven True Graph-driven $\bm{A}$.
infer_x_c_together False Infer $\bm{x}$ and $\bm{c}$ together.
include_mask True Include mask.
norm_D_cols False Normalize columns of $\bm{A}$.
lambda_D 0.3 Regularization parameter for $\bm{A}$.
step_D_decay 0.9999 Decay rate for step size of $\bm{A}$.
l1_D 2.74 $\ell_{1}$ regularization for $\bm{A}$.
Appendix F Information about the baselines:
  1. 1.
    Without Dynamic Prior”: Running CrEIMBO while removing the prior over the dynamics from the inference. Instead, inferring the temporal traces of the ensembles using regularized least-squares with the addition of smoothness, de-correlation, and Frobenius norm regularization terms on the dynamics.
  2. 2.
    All Regions (sparse)”: Running CrEIMBO without applying the multi-regional diagonal mask over the ensemble compositions ($\bm{A}$), thus supporting the finding of non-localized ensembles.
  3. 3.
    All Regions (sparse)”: Similar to the previous one but without applying sparsity on the ensemble matrix.
  4. 4.
    PCA All Regions”: Running CrEIMBO while replacing the dimensionality reduction step with PCA.
  5. 5.
    Single Session #”: Running CrEIMBO on a single session (view) of the data rather than leveraging cross-session information.
    Since $\{\bm{f}_{k}\}$ is invariant to its ordering, we used SciPy’s implementation of the ‘linear_sum_assignment‘ problem [4] to re-order the identified sub-circuits before comparing them with the ground truth, matching the $f$ values by minimizing the $\ell_{2}$ error.
Appendix G Code and Data Sharing
The neural data used in this study [12] is publicly available with License: “spdx:CC-BY-4.0”. We used version 0.240123.1806 of the data which can be accessed through the DANDI Archive at https://dandiarchive.org/dandiset/000469/0.240123.1806 (Access Information: dandi:OpenAccess). The code will be shared via GitHub upon publication.
Appendix H Technical Details
The whole code is written in Python. We utilized the following Python packages along with their respective versions: scipy (1.8.0), matplotlib (3.8.2), numpy (1.23.5), pandas (1.5.0), seaborn (0.11.2), scikit-learn (1.0.2), pylops (1.18.2), and statsmodels (0.13.2), all within a Python 3.10.4 environment. For the $\ell_{1}$ regularization, we used the “SPGL1” solver of pylops [27]. During the writing process, we used ChatGPT very sparingly to improve wording and fix grammar mistakes at the word or sentence level. The synthetic experiment can run locally on a desktop (simple ASUS machine). However, the robustness tests with hundreds of random repeats and the human data experiments used the entry node with CPUs only of NERSC’s “Perlmutter” supercomputer. Running time for each synthetic instance ranges from 0.8 to 2 minutes (with the capability to run in parallel). For the human data, processing all 41 sessions with 1000 time points per session (and varying neuron counts as described in the text and figures) takes between 17 to 25 minutes, depending on parameters and result quality.
Appendix I Societal Impact and Potential Harmful Consequences
With respect to Societal Impact, the development, application, and future work of CrEIMBO can advance our understanding of brain function. Future work can potentially lead to improved diagnostics of neurological disorders. We believe our method does not present any significant harmful consequences. We explicitly discussed the method’s limitations, ensured proper use of the data, and adhered to the NeurIPS code of ethics.

### v2

@@
-CrEIMBO: Cross Ensemble Interactions in Multi-view Brain Observations
+CREIMBO: Cross-Regional Ensemble Interactions in Multi-view Brain Observations
 Abstract
-Modern recordings of neural activity provide diverse observations of neurons across brain areas, behavioral conditions, and subjects—thus presenting an exciting opportunity to reveal the fundamentals of brain-wide dynamics underlying cognitive function.
-Current analysis methods, however, often fail to fully harness the richness of such data as they either provide an uninterpretable representation (e.g., via “black box” deep networks) or over-simplify the model (e.g., assume stationary dynamics or analyze each session independently).
-Here, instead of regarding asynchronous recordings that lack alignment in neural identity or brain areas as a limitation, we exploit these diverse views of the same brain system to learn a unified model of brain dynamics.
-Specifically, we assume that brain observations stem from the joint activity of a set of functional neural ensembles (groups of co-active neurons) that are similar in functionality across recordings, and propose to discover the ensemble structures and their non-stationary dynamical interactions in a new model we term CrEIMBO (Cross-Ensemble Interactions in Multi-view Brain Observations).
-CrEIMBO identifies the composition of the per-session neural ensembles through graph-driven dictionary learning and models the ensemble dynamics as a latent sparse time-varying decomposition of global sub-circuits—thereby capturing non-stationary dynamics.
-CrEIMBO identifies multiple co-active sub-circuits while maintaining representation interpretability due to sharing sub-circuits across sessions.
-Furthermore, CrEIMBO distinguishes session-specific from global (session-invariant) computations by exploring when distinct sub-circuits are active.
-We demonstrate CrEIMBO’s ability to recover ground truth components in synthetic data and uncover meaningful brain dynamics, capturing cross-subject and inter- and intra-area variability, in high-density electrode recordings of humans performing a memory task.
-Figure 1: Challenges and approach.
-Top right: Representation of the latent $K$ neural ensemble sub-circuits ($\{\bm{f}_{k}\}$) whose time-changing contributions to the full dynamics ($\bm{F}_{t}=\sum_{k=1}^{K}\bm{c}_{kt}\bm{f}_{k}$) governs the whole-brain activity.
-These subcircuits can either capture global task-related patterns (e.g.
-$\bm{f}_{1}$), capture general brain computations that vary over time and between sessions (e.g.
-$\bm{f}_{2}$), be active only under a specific session (e.g.
-$\bm{f}_{3}$), or capture both task related and general per-session computations at different period ($\bm{f}_{k}$, e.g., processing related to some visual cue presented at the beginning of the session only ).
-Top Left: Membership of neurons within functional neural ensembles.
-Bottom: Real-world data often provide distinct partial views (“sessions”) of the brain system that cannot be matched in terms of neural identity, quantity, or distribution between areas.
-This discrepancy may lead to biased conclusions, e.g., session 2 and session D lack representations from areas $a$ and $d$, respectively.
+Modern recordings of neural activity provide diverse observations of neurons across brain areas, behavioral conditions, and subjects; presenting an exciting opportunity to reveal the fundamentals of brain-wide dynamics.
+Current analysis methods, however, often fail to fully harness the richness of such data, as they provide either uninterpretable representations (e.g., via deep networks) or oversimplify models (e.g., by assuming stationary dynamics or analyzing each session independently).
+Here, instead of regarding asynchronous neural recordings that lack alignment in neural identity or brain areas as a limitation, we leverage these diverse views into the brain to learn a unified model of neural dynamics.
+Specifically, we assume that brain activity is driven by multiple hidden global sub-circuits.
+These sub-circuits represent global basis interactions between neural ensembles—functional groups of neurons—such that the time-varying decomposition of these sub-circuits defines how the ensembles’ interactions evolve over time non-stationarily and non-linearly.
+We discover the neural ensembles underlying non-simultaneous observations, along with their non-stationary evolving interactions, with our new model, CREIMBO (Cross-Regional Ensemble Interactions in Multi-view Brain Observations).
+CREIMBO identifies the hidden composition of per-session neural ensembles through novel graph-driven dictionary learning and models the ensemble dynamics on a low-dimensional manifold spanned by a sparse time-varying composition of the global sub-circuits.
+Thus, CREIMBO disentangles overlapping temporal neural processes while preserving interpretability due to the use of a shared underlying sub-circuit basis.
+Moreover, CREIMBO distinguishes session-specific computations from global (session-invariant) ones by identifying session covariates and variations in sub-circuit activations.
+We demonstrate CREIMBO’s ability to recover true components in synthetic data, and uncover meaningful brain dynamics in human high-density electrode recordings, including cross-subject neural mechanisms as well as inter- vs.
+intra-region dynamical motifs.
+Furthermore, using mouse whole-brain recordings, we show CREIMBO’s ability to discover dynamical interactions that capture task and behavioral variables and meaningfully align with the biological importance of the brain areas they represent.
+Figure 1: CREIMBO’s Illustration.
+A: Real-world multi-regional neural datasets consist of multiple ($D$) non-simultaneous recording sessions ($\{{\bm{Y}^{d}}\}^{D}_{d=1}$) that cannot be matched in terms of individual neurons’ identities, quantity, or cross-region distribution, resulting in neurons appearing only in certain sessions (4-pointed stars) or not at all (gray circles).
+This variability hinder our ability to draw unified conclusions of whole-brain neural computation, while analyzing sessions individually could lead to session-specific biased results.
+This challenge is further complicated by variability in session (or trial) duration, the non-stationarity of brain dynamics (even within a session), and the presence of multiple trials within each session.
+B $\rightarrow$ A: Instead of considering these asynchronous sessions as a challenge, we frame them as an opportunity to obtain distinct, potentially complementary “views” onto the shared underlying brain system, thereby facilitating the learning of a unified brain dynamics model.
+We assume that brain computations are mediated by multiple neural ensembles—$\bigstar$ groups of same-region neurons with shared functionalities—whose interactions yield meaningful neural representations.
+However, the specific neurons and their membership degrees within each ensemble are unknown, and a neuron may belong to multiple ensembles with varying membership levels based on its diverse functionalities.
+C: CREIMBO leverages partial brain views to co-learn ensemble compositions per session ($\{\bm{A}^{d}\}_{d=1}^{D}$), their temporal activity ($\{\bm{x}^{d}\}_{d=1}^{D}$), and cross-regional interactions $\bm{x}^{d}_{t+1}=\bm{F}_{t}^{d}\bm{x}^{d}_{t}$.
+It posits that these non-stationary interactions stem from a session-shared dictionary prior of up to $K$ global interactions $\{\bm{f}_{k}\}_{k=1}^{K}$, whose sparse time-changing decomposition $\bm{F}^{d}_{t}=\sum_{k=1}^{K}\bm{f}_{k}\bm{c}_{kt}^{d}$ shapes the overall ensemble dynamics at each time point $t$ and session $d$.
+The interactions’ coefficients can help distinguish session-specific interactions from session-invariant interactions in different time periods (pink right, solid vs.
+dotted in subplot C right).
+With the session-shared ensemble-interactions dictionary and certain assumptions (Sec.
+Assumptions), the model auto-sorts ensembles across sessions by functionality, aligning the $j$-th ensemble of session $d$ with the $j$-th ensemble of session $d^{\prime}$, despite differences in observed neurons.
 1 Introduction
@@
 Individual recording sessions provide a singular perspective of the brain dynamics, as the reinsertion of recording devices results in the capture of different subsets of neurons in each session.
-Thus fitting population-level models often reverts to per-session analysis (i.e., fitting each session independently), which is inherently constrained by the inability to incorporate the full set of recorded activity across all recorded brain areas across all sessions.
-Moreover, analyzing the data session-by-session, may hinder our ability to distinguish between computations that are session or subject specific from task related or global session invariant neural processes.
+Thus, fitting population-level models often reverts to per-session analysis (i.e., fitting each session independently), which is inherently constrained by the inability to incorporate the full set of recorded activity across all recorded brain areas across all sessions.
+Moreover, analyzing the data session-by-session, may hinder our ability to distinguish between computations that are session or subject specific from task related or global session-invariant neural processes.
 Compounding the difficulty of merging different sessions’ data are the non-linear and non-stationary nature of neural activity, as well as the presence of noise and variability across trials—both of which require additional model complexity to account for.
-Moreover to extract scientifically meaningful insight, the models must be interpretable in that model parameters must be readily relatable to task variables or connections between recorded units.
+Moreover, to extract scientifically meaningful insights, the models must be interpretable, with model parameters directly related to task variables or connections between recorded units.
 This gap between the opportunities offered by modern neural data and the limited capabilities of current methods necessitates new approaches to leverage the richness in modern brain data and discover the fundamental neural sub-circuits governing brain activity.
-One current approach used to combine multi-session data into a single model is the use of deep learning models.
-By training to predict all sessions’ data, deep learning models implicitly merge the datasets into the fit network weights.
-While these powerful models can combine information across sessions, their uninterpretable ‘black box’ nature and complexity typically hinders their ability to reveal the fundamental building blocks of neural computation.
-Other approaches, including tensor factorization and its variations (e.g., [3, 9]) may be limited in their ability to merge sessions of varying durations or identify the interactions between the identified neural components.
-Here, we re-frame cross-session variability in neurons recorded and experimental settings not as a drawback, but as a valuable advantage.
-Rather than viewing the different sessions as capturing distinct neural populations, we perceive such data as providing multiple, complementary perspectives of one brain system.
+One current approach used to combine multi-session data into unified model is deep learning.
+By training to predict all sessions’ data, deep networks implicitly merge the datasets into the fit network weights.
+While these powerful models can combine information across sessions, their uninterpretable nature and complexity typically hinders their ability to reveal the fundamental building blocks of neural computation.
+Other approaches, including tensor factorization and its variations (e.g., Chen et al.
+(2015); Harshman (1970)) may be limited in their ability to merge sessions of varying durations or model the interactions between the identified neural components.
+Here, we re-frame cross-session variability in neurons recorded and experimental settings, not as a drawback, but as a valuable advantage, viewing this data as providing multiple, complementary “views” into a single, shared brain system (Fig.
+1B $\rightarrow$ A).
 This approach enables a more thorough and holistic discovery of the underlying system designed to extract joint information from the entire dataset.
-In particular, we hypothesize that the population dynamics at each session is driven by a common set of global sub-circuit interactions shared across sessions.
-To capture the non-linear and non-stationary aspects of the dynamics, we leverage recent advances in decomposed linear dynamical systems (dLDS), which enables us to model how the neural trajectories transverse a low-dimensional manifold via a time-varying combination of the sub-circuits.
-While we assume that the representation of the latent dynamics is shared across sessions, the measured neural ensembles are not.
-We thus model each session’s ensembles as independent neural representations, i.e., the membership of neurons to ensembles, that captures how the latent trajectories project into each sessions’ view of the brain.
-In this work we lay out this novel model, which we term Cross Ensemble Interactions in Multi-view Brain Observations (CrEIMBO), and demonstrate its ability to capture latent dynamics in multi-view neural data.
-Specifically our contributions include:
+In particular, we hypothesize that the population dynamics at each session is driven by a common set of global time-invariant sub-circuits shared across sessions.
+To effectively capture the non-linear and non-stationary dynamics of the brain, we enable these sub-circuits to adjust their activity levels over time and across sessions.
+This approach allows us to identify evolving neural patterns through the collective activity of the sub-circuits, modeled via their time-varying decomposition.
+We assume that the latent dynamics governed by functional neural groups are shared across sessions, even though the specific individual neurons observed may vary.
+To address this, we propose a per-session transformation from the joint low-dimensional circuit space to each session’s observation space, ensuring cross-session alignment of the neural groups’ functionalities.
+In this work, we lay out this new model, which we term Cross-Regional Ensemble Interactions in Multi-view Brain Observations (CREIMBO), and demonstrate its ability to capture latent dynamics in multi-view neural data.
+Specifically, our contributions include:
   • We discover multi-regional brain dynamics through leveraging the richness of modern neural datasets while ensuring interpretability.
-  • We distinguish between intra- and inter-region interactions by identifying sparse, same-area neural ensembles, deploying a specialized structured prior over per-session projections from circuit space to neural space.
-  • We accurately recover ground truth components in synthetic data and discover meaningful cross-regional brain interactions underlying human brain data utilizing a shared low-dimensional circuit space across recording sessions.
+  • We distinguish between intra- and inter-region interactions by identifying sparse, same-area neural ensembles, deploying a specialized structured prior over per-session projections from the circuit latent space to the observed neural space.
+  • We accurately recover ground truth components in synthetic data and discover meaningful cross-regional brain interactions underlying human and mouse brain data.
 2 Related Work
@@
 Merging non-simultaneous neural recording sessions.
-Existing work for combining non-simultaneous neural population recordings, here termed “sessions”, include  [26] who introduced the term “stitching” for this problem and modeled neural activity with a shared latent dynamical systems model and [24] who proposed to stitch asynchronous recordings for inferring functional connectivity by assuming shared cross-session latent input and demonstrated the advantage in observing more neurons for inferring connectivity.
-However, none of these methods capture multiple ($>2$) interacting neural sub-circuits active simultaneously spanning populations across multiple areas, nor do they distinguish between within- and between-area interactions, nor do they address temporal non-stationarity.
-In [1, 2], Bishop et al.
-proposed integrating the neural activity structure into inference and expanding the proposed applications of stitching to study of communication between two neural populations of brain areas and proposed a theoretical framework of when stitching is possible.
-In [20], the authors proposed extracting low-dimensional dynamics from multiple sessions by learning temporal co-variations across neurons and time.
+Existing work for combining non-simultaneous neuronal population recordings, here termed “sessions”, include  Turaga et al.
+(2013) who introduced the term “stitching” for this problem and modeled neural activity with a shared latent dynamical systems model and Soudry et al.
+(2013) who proposed to stitch asynchronous recordings for inferring functional connectivity by assuming shared cross-session latent input and demonstrated the advantage in observing more neurons for inferring connectivity.
+However, these methods are not designed to capture more than two simultaneously interacting neural sub-circuits across multiple populations, nor do they distinguish between within- and between-area interactions or specifically address temporal non-stationarity.
+Bishop (2015) and Bishop & Yu (2014) proposed integrating the neural activity structure into inference and expanding stitching to study the communication between two neuronal populations from potentially different brain areas.
+In Nonnenmacher et al.
+(2017), the authors proposed extracting low-dimensional dynamics from multiple sessions by learning temporal co-variations across neurons and time.
 Both of these methods, through, assume at least a small overlap in the identity of cross-session observed neurons.
-LFADS [11, 25] enables stitching while addressing the need for neural overlap by finding a shared non-linear dynamical model.
-However, the latent dynamics may be uninterpretable in terms of interactions among neural populations, and due to its foundation on a variational auto-encoder, it may be challenging to tune.
+LFADS (Keshtkaran et al., 2021; Pandarinath et al., 2018) enables stitching while addressing the need for limited neural overlap by finding a shared non-linear latent dynamical model.
+However, its latent dynamics may be difficult to interpret regarding the specific neural sub-circuits that compose it, and due to its foundation on a variational auto-encoder, it may be challenging to tune.
 Maintaining Interpretability in Non-Stationary Neural Dynamics Models.
 Real-world neural activity follows non-stationary nature due to changing environmental settings.
-Classical dynamical models often capture non-stationarity through various methods such as ’black box’ deep learning models (e.g.
-[22, 25, 29], which are powerful but are often not interpretable with respect to neural interactions; linear models, which are interpretable but not expressive enough to capture the complex nature of neural data; or switching piece-wise linear models  [15, 14, 19], which fail to enable the identification of co-active dynamics.
-dLDS (decomposed Linear Dynamical Systems) [16] (with its improved learning procedure in [18]), addresses this gap by generalizing the switching assumption to sparse time-changing decomposition of linear dynamical systems, thus enabling capturing co-active sub-circuits.
-However, while the current dLDS framework can be modified for co-processing multi-subject recordings that present variability in neuron identity and distribution, this modification lacks the structured regional prior over the loading matrix that is necessary to partition multi-area recordings such that between and within area interactions can be isolated and studied.
-Identification of Functional Neural Ensembles A potential way to find interpretable loading matrix that links low to high dimensional space is through the identification of functional neural ensembles—here refereed to as groups of neurons with similar functionalities—that determine the axes in the latent space.
+Classical dynamical models often simplified neural dynamics to linear (e.g., jPCA Churchland et al.
+(2012)) or non-linear but stationary (e.g., Gallego et al.
+(2020)).
+Models that capture non-stationarity often derive from “black box” deep learning models (e.g., Schneider et al.
+(2023); Pandarinath et al.
+(2018); Zhu et al.
+(2022), which are powerful but often present limited interpretability with respect to neural interactions.
+Other models are built on switching piece-wise linear models  Linderman et al.
+(2016; 2017); Murphy (1998), which do not enable the identification of multiple co-active dynamic processes.
+dLDS (decomposed Linear Dynamical Systems,  Mudrik et al.
+(2024a)) and its multi-step extension (Mudrik et al., 2024c)), addresses this gap by generalizing the switching assumption to sparse time-changing decomposition of linear dynamical systems, thus enabling capturing co-active sub-circuits.
+However, while dLDS can be extended to co-process multi-session recordings with cross-session variability in neuron identity, this modification may not yield interpretable latent-to-observed projections and lacks the structured multi-regional prior over the emission matrix necessary to partition multi-area recordings and study between- vs.
+within-area interactions.
+Identification of Functional Neural Ensembles.
+A potential way to find interpretable loading matrix that links low to high dimensional space is through the identification of functional neural ensembles—here referred to as groups of neurons with similar functionalities—that determine the axes in the latent space.
 For instance, the mDLAG model used Automatic Relevance Determination (ARD) to promote population-wise sparsity patterns of each latent variable.
-In [17], the authors presented an approach to address sparsity in loading matrices while accounting for multi-trial and task-conditions variability.
+In Mudrik et al.
+(2024b), the authors presented an approach to address sparsity in loading matrices while accounting for multi-trial and task-conditions variability.
 They proposed a graph-driven regularization technique that facilitates the grouping of co-active neurons into the same ensembles and pushing apart neurons that present different functionality, while allowing for structural adjustments in ensemble configurations across different task conditions.
@@
 However, their model does not model the dynamic interactions between ensembles or give a closed form prior for the evolution of the ensembles’ temporal traces.
-While other methods, e.g., clustering approaches like  [8] or Sparse Principal Component Analysis (SPCA) [30], enable the recognition of underlying sparse groups, they either do not allow a neuron to belong to multiple groups with varying degrees of membership or do not support controlled adaptations of structural membership across sessions.
+While other methods, e.g., clustering approaches like  (Grossberger et al., 2018) or Sparse Principal Component Analysis (SPCA) Zou et al.
+(2006), enable the recognition of underlying sparse groups, they either do not allow a neuron to belong to multiple groups with varying degrees of membership or do not support controlled adaptations of structural membership across conditions.
 Multi-regional brain interactions.
-With the ability to record from multiple brain areas and the recognition of the importance of considering multi-area communication [21], recent works have delved into understanding the communication between or within brain areas.
-In [7, 6], the authors proposed models to identify time delays between cross-regional brain interactions, overcoming the challenge of distinguishing indirect, concurrent, and bi-directional interactions in two or more populations.
+With the ability to record from multiple brain areas and the recognition of the importance of considering multi-area communication (Pesaran et al., 2021), recent works have delved into understanding the communication between or within brain areas.
+In Gokcen et al.
+(2022; 2024), the authors proposed models to identify time delays between cross-regional brain interactions, overcoming the challenge of distinguishing indirect, concurrent, and bi-directional interactions in two or more populations.
 However, their model focuses on analyzing sessions individually and is not intended to uncover the full underlying set of transition matrices, but rather to recognize meaningful time delays that can imply interactions.
-Other models addressed multi-regional interactions through a communication subspace [23] using dimensionality reduction, or by Generalized Linear Models (GLMs), with either Poisson [28] or Gaussian [28] statistics to identify functional coupling between areas.
+Other models addressed multi-regional interactions through a communication subspace Semedo et al.
+(2019) using dimensionality reduction, or by Generalized Linear Models (GLMs), with either Poisson (Yates et al., 2017) or Gaussian (Yates et al., 2017) statistics to identify functional coupling between areas.
 These approaches, though, are not tailored to capture non-simultaneous cross-sessions variability in neuron identity and subject processing.
-Other recent methods, including [10, 5, 13] discussed the potential usage of their model to multi-regional interactions, have proposed studying multi-regional interactions using switching dynamical systems approaches, where the model switches between different states over time to capture distinct periods.
-However, this approach cannot disentangle interactions among multiple sub-circuits active simultaneously, nor can it identify interactions at multiple temporal resolutions that occur concurrently and distinguish between these or multiple co-processed processes whose encoding occurs gradually rather than through sharp transitions.
-3 Problem Formulation and Approach
-Problem Formulation
-Let $\{\bm{Y}^{d}\}_{d=1}^{D}$ be a set of estimated firing rates over $D$ asynchronous sessions, indexed by ${d=1\dots D}$.
-The neural recordings of each $d$ session, $\bm{Y}^{d}\in\mathbb{R}^{N^{d}\times T^{d}}$ captures $T^{d}$ time observations of the activity of $N^{d}$ neurons from up to $J$ distinct brain areas, thus offering a partial view of the brain system.
+Other recent methods (Karniol-Tambour et al.
+(2024); Glaser et al.
+(2020); Li et al.
+(2024)) discussed the potential usage of their model to multi-regional interactions, have proposed studying multi-regional interactions using switching dynamical systems approaches, where the model switches between different states over time to capture distinct periods.
+However, this approach cannot disentangle interactions among multiple sub-circuits active simultaneously, nor can it identify interactions at multiple temporal resolutions that occur concurrently and distinguish between these or multiple co-processed processes whose encoding occurs gradually rather than via sharp transitions.
+3 Problem Introduction and Approach
+Problem Intuition: Let $\{\bm{Y}^{d}\}_{d=1}^{D}$ be a set of estimated neural firing rates (e.g.
+from neuropixels data) over $D$ asynchronous sessions, indexed by ${d=1\dots D}$.
+The neural recordings of each $d$ session, $\bm{Y}^{d}\in\mathbb{R}^{N^{d}\times T^{d}}$, capture $T^{d}$ time observations of the activity of $N^{d}$ neurons from up to $J$ distinct brain areas, thus offering a partial view of the brain system.
 Note that sessions may refer to the same or different subjects, can vary in duration or number of neurons, i.e.
-$N^{d}$ and $T^{d}$ need not be equal to $N^{d^{\prime}}$ and $T^{d^{\prime}}$ for different sessions $d,d^{\prime}$, and may differ in the sub-set of brain areas (out of $J$) they capture.
+$N^{d}$ and $T^{d}$ need not be equal to $N^{d^{\prime}}$ and $T^{d^{\prime}}$ for different sessions $d,d^{\prime}$, and can differ in the subset of brain areas (out of $J$) they capture.
 Hence, these cross-session data matrices cannot be aligned into a single data array, e.g., via a tensor—hindering the direct application of existing analyses as tensor factorization.
@@
 Moreover, these sub-circuits may be active simultaneously or intermittently, yielding a model that can flexibly fit neural trajectories that differ between sessions with a single underlying mechanism.
-The problem we address in this model is identifying the set of unknown multi-regional sub-circuits, including their neuronal composition and non-stationary activation levels, from asynchronously collected observations.
-CrEIMBO (Cross Ensemble Interactions in Multi-view Brain Observations)
-A naive approach to identifying the unknown sub-circuits involves first projecting each session’s data into a low-dimensional space and subsequently fitting the dynamics model in the low-dimensional space.
+The problem CREIMBO addresses is thus identifying the set of $K$ latent multi-regional sub-circuits $\{\bm{f}_{k}\}_{k=1}^{K}$, capturing interactions between unknown functional sparse neural ensembles $\bm{A}$, along with their non-stationary activation levels $\{\{\bm{c_{kt}^{d}}\}_{k=1}^{K}\}_{d=1}^{D}$, by leveraging the joint information from asynchronously collected observations $\{\bm{Y}^{d}\}_{d=1}^{D}$.
+CREIMBO: A naive approach to identifying the unknown sub-circuits involves first projecting each session’s data into a low-dimensional space and subsequently fitting the dynamics model in the low-dimensional space.
 However, this two-step approach can be sub-optimal as the dimensionality reduction projections may not prioritize the same dimensions needed to preserve dynamic fidelity.
@@
 Let $\bm{A}^{d}$ be a sparse matrix where each column, $\bm{A}^{d}[:,j]$, encodes the neuronal composition of the $j$-th ensemble in the $d$-th session, such that non-zero values within these columns represent the membership of neurons in the ensemble and their magnitudes reflecting the degree of membership.
-Since we aim to capture multi-regional interactions, we design the ensemble matrix $\bm{a}:=\bm{A}^{d}$ as a block diagonal matrix where the $j^{th}$ block, $\bm{a}^{j}\in\mathbb{R}^{n_{j}\times p_{j}}$, contains the neural ensembles for the $j$-th area.
+Since we aim to capture multi-regional interactions, we design the ensemble matrix $\bm{a}:=\bm{A}^{d}$ as a block diagonal matrix where the $j$-th block, $\bm{a}^{j}\in\mathbb{R}^{n_{j}\times p_{j}}$, contains the neural ensembles for the $j$-th area.
 We denote $p_{j}$ as the number of ensembles within that area (note that $p=\sum_{j=1}^{J}p_{j}$) and $n_{j}:=N_{j}^{d}$ as the number of neurons in that area (i.e., the total number of neurons observed in session $d$ is $N^{d}=\sum_{j=1}^{J}n_{j}$).
-Furthermore, let $\tau:=\mathrm{T}_{m}^{d}$ indicate the time duration (i.e., the number of time-points) of the $m^{th}$ trial, for $m=1\dots M_{d}$, in session $d$.
-In each trial, the temporal activity for the $p$ ensembles are denoted as $\bm{x}:=\bm{X}_{m}^{d}\in\mathbb{R}^{p\times\tau}$, leading to a model where the observations $\bm{y}:=\bm{Y}_{m}^{d}$ in trial $m$ are assumed to arise from the joint activity of all these ensembles, up do a normally-distributed error $\epsilon\sim\mathcal{N}(0,\sigma)$, i.e., $\bm{y}=\bm{a}\bm{x}+\epsilon.$ The evolution of the ensembles’ activities ($\bm{x}$) reflect the latent interactions between ensembles, which we model with a non-stationaly linear dynamical system $\bm{x}_{t}=\bm{F}_{t}\bm{x}_{t-1}$ for $t=1\ldots\tau$.
+Furthermore, let $\tau:=\mathrm{T}_{m}^{d}$ indicate the number of time-points recorded in each $m$-th trial of session $d$ ($m=1\dots M_{d}$), such that the observations in that trial are captured by $\bm{Y}_{m}^{d}\in\mathbb{R}^{N^{d}\times\mathrm{T}_{m}^{d}}$.
+In each trial $m$, the temporal activity for the $p$ ensembles are denoted as $\bm{x}:=\bm{X}_{m}^{d}\in\mathbb{R}^{p\times\tau}$, leading to a model where the observations $\bm{y}:=\bm{Y}_{m}^{d}$ in trial $m$ are assumed to arise from the joint activity of all these ensembles, up do a normally-distributed error $\epsilon\sim\mathcal{N}(0,\sigma)$, i.e., $\bm{y}=\bm{a}\bm{x}+\epsilon.$ Please refer to Section G for Poisson statistics inference in low spiking rate regimes.
+The evolution of the ensembles’ activities ($\bm{x}$) reflect the latent interactions between ensembles, which we model with a non-stationary linear dynamical system $\bm{x}_{t}=\bm{F}_{t}\bm{x}_{t-1}$ for $t=1\ldots\tau$.
 The time-changing transition matrix $\bm{F}_{t}\in\mathbb{R}^{p\times p}$ captures both non-linear and non-stationary brain activity.
 A key desired property of our model is to capture the overlapping activity of multiple sub-circuits of interacting ensembles.
-We thus follow the dLDS concept and model the interactions ($\bm{F}_{t}$) through a time-varying sparse decomposition ${\bm{F}_{t}=\sum_{k=1}^{K}\bm{c}_{kt}\bm{f}_{k}}$, where $\{\bm{f}_{k}\}_{k=1}^{K}$ represents the $K$ global ensemble interactions, and $\bm{c}_{kt}$ are the time-varying coefficients that represent the modulation of each interaction pattern over time.
-Fitting this model requires identifying the global operators $\{\bm{f}_{k}\}_{k=1}^{K}$ and the functional ensembles $\{\bm{A}^{d}\}_{d=1}^{D}$.
-For this, we employ an alternating approach, where we iterate between updating the global operators, $\{\bm{f}_{k}\}_{k=1}^{K}$, and infer, for each session $d$, 1) the ensemble compositions $\bm{a}:=\bm{A}^{d}$, 2) their dynamics $\bm{x}$, and 3) the sub-circuits’ temporal coefficients $\bm{c}_{t}$.
-These steps are iterated until the parameters converge.
+We thus follow dLDS and model the interactions ($\bm{F}_{t}$) through a time-varying sparse decomposition ${\bm{F}_{t}=\sum_{k=1}^{K}\bm{c}_{kt}\bm{f}_{k}}$, where $\{\bm{f}_{k}\}_{k=1}^{K}$ are the $K$ global ensemble interactions, and $\bm{c}_{kt}$ are their time-varying coefficients that capture the modulation of each interaction.
+Fitting this model requires identifying the global operators $\{\bm{f}_{k}\}_{k=1}^{K}$ and ensembles $\{\bm{A}^{d}\}_{d=1}^{D}$.
+For this, we employ an alternating approach and iterate between updating $\{\bm{f}_{k}\}_{k=1}^{K}$ and inferring, for each session $d$, 1) the ensemble compositions $\bm{a}:=\bm{A}^{d}$, 2) their dynamics $\bm{x}$, and 3) the sub-circuits’ temporal coefficients $\bm{c}_{t}$.
+These steps are iterated until convergance.
 Ensemble update: We update $\bm{a}$ per-row (neuron) $n$ ($\bm{a}_{n}$).
-Following the work of [17], we infer the sparse ensemble structures with graph-driven re-weighted $\ell_{1}$ regularization that groups together neurons with shared activity patterns, while pushing apart neurons who do not behave similarly.
+Following the work of Mudrik et al.
+(2024b), we infer the sparse ensemble structures with graph-driven re-weighted $\ell_{1}$ regularization that groups together neurons with shared activity patterns, while pushing apart neurons who do not behave similarly.
 The graph for each area $j$, $\bm{h}^{j}\in\mathbb{R}^{n_{j}\times n_{j}}$, is calculated based on a data-driven Gaussian kernel that measures the temporal similarity between neurons, such that $\bm{h}^{j}_{n1,n2}=exp(\frac{\|\bm{y}^{j}_{n1}-\bm{y}^{j}_{n2}\|_{2}^{2}}{\sigma_{h}})$ where $\sigma_{h}$ is a hyperparameter.
 Specifically, for each row $\bm{a}_{n,:}$ and considering only the ensembles in region $j_{n}$ that neuron $n$ is part of,
-$\widehat{\bm{a}}_{n,:}=\arg\min_{\bm{a}_{n:}}\|\bm{y}_{n:}-\bm{a}_{n:}\bm{x}_{j_{n}:}\|_{2}^{2}+\sum_{j=1}^{p}\bm{\lambda}^{d}_{n,j}|\bm{a}_{n,j}|\quad\textrm{ where }\quad\bm{\lambda}_{n,j}=\frac{\beta_{1}}{\beta_{2}+|\widehat{\bm{a}}_{n,j}|+\beta_{3}|\bm{h}_{n:}^{j}\widehat{\bm{a}}_{:j}|}.$
-Here, $\bm{x}_{j_{n},:}$ is the activity of the ensembles of the $j_{n}^{th}$ region and $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are scalars that control the effect of the graph on the regularization.
+$\widehat{\bm{a}}_{n,:}=\arg\min_{\bm{a}_{n:}}\|\bm{y}_{n:}-\bm{a}_{n:}\bm{x}_{j_{n}:}\|_{2}^{2}+\sum_{j=1}^{p}\bm{\lambda}^{d}_{n,j}|\bm{a}_{n,j}|\textrm{ where }\bm{\lambda}_{n,j}=\frac{\beta_{1}}{\beta_{2}+|\widehat{\bm{a}}_{n,j}|+\beta_{3}|\bm{h}_{n:}^{j}\widehat{\bm{a}}_{:j}|}.\vspace{-5pt}$ (1)
+Here, $\bm{x}_{j_{n},:}$ is the activity of the ensembles of the $j_{n}$-th region and $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are scalars that control the effect of the graph on the regularization.
 The graph adjacency $\bm{h}_{n:}\widehat{\bm{a}}_{:j}$ weights the sparse membership in the $j$-th ensemble based on temporal similarity.
@@
 Latent state and dynamics coefficient update: The ensemble activities $\bm{x}$ and coefficients $\bm{c}$ are updated iterative for each time-point $t$ via the LASSO optimization
-${\widehat{\bm{x}}_{t},\widehat{\bm{c}}_{t}=\arg\min_{\bm{x}_{t},\bm{c}_{t}}\|\bm{y}_{t}-\bm{a}\bm{x}_{t}\|_{2}^{2}+\left\|\bm{x}_{t}-\left(\sum_{k=1}^{K}\bm{f}_{k}\bm{c}_{k,(t-1)}\right)\bm{x}_{t-1}\right\|_{2}^{2}+\lambda_{x}\|\bm{x}_{t}\|_{1}+\lambda_{c}\|\bm{c}_{t}\|_{1}}$ (1)
-where $\lambda_{c}\in\mathbb{R}$ and $\lambda_{x}\in\mathbb{R}$ are regularization weights.
+${\widehat{\bm{x}}_{t},\widehat{\bm{c}}_{t}=\arg\min_{\bm{x}_{t},\bm{c}_{t}}\|\bm{y}_{t}-\bm{a}\bm{x}_{t}\|_{2}^{2}+\left\|\bm{x}_{t}-\left(\sum_{k=1}^{K}\bm{f}_{k}\bm{c}_{k,(t-1)}\right)\bm{x}_{t-1}\right\|_{2}^{2}+\lambda_{c}\|\bm{c}_{t}\|_{1}}$ (2)
+where $\lambda_{c}\in\mathbb{R}$ is sparsity-regularization weights on the basis interactions’ activations.
 Dynamical system update: The global ensembles interactions, $\{\bm{f}_{k}\}_{k=1}^{K}$ are assumed to be sparse (i.e. ensembles interactions are not all to all) and are identified directly by:
-$\bm{F}^{\textrm{all}}_{n}=\arg\min_{{\bm{F}^{\textrm{all}}}_{j}}\|\bm{y}_{n}-\bm{F}_{n}^{\textrm{all}}\bm{cx}\|+\lambda_{F}\|\bm{F}^{\textrm{all}}_{j}\|_{1}+\sum_{{k_{1}},{k_{2}},({k_{1}}\neq{k_{2}})}\rho({\bm{f}}_{k_{1}},{\bm{f}}_{k_{2}})$ (2)
-where $\bm{F}^{\textrm{all}}$ is an horizontal concatenation of all $\bm{f}$s, and $\sum_{{k_{1}},{k_{2}}}\rho(\bm{f}_{k_{1}},\bm{f}_{k_{2}})$ is a de-correlation term used to ensure that distinct $\bm{f}$s are not too similar, Above, $\bm{cx}_{t}:=(\bm{c}^{batch}_{t}\otimes[1]_{1\times p})\circ([1]_{1\times K}\otimes\bm{x}^{batch}_{t})$, and $\lambda_{F}$ is the weight on the sparsity-regularization on $\bm{F}$ and “batch” means that this operation is in practice taken on a random subset of sessions for computational complexity considerations.
-Figure 2: A-E: Testing CrEIMBO over 312 iterations of random initializations and parameters (see Sec.
-D) reveals high correlations between ground truth and fit components for all unknowns, $\bm{Y}$, $\bm{A}$, $\bm{x}$, $\bm{F}$, $\bm{c}$ (subplots A-E respectively).
-F-G: A comparison between the ensemble trajectory ($\bm{X}$) and the ensemble compositions ($\bm{A}$) as identified by CrEIMBO and other approaches compared to the ground-truth.
-Details about the other approaches are found in Section F.
-Additional sessions in Figure 5.
-H-J: $-\log(MSE)$ between the ground truth latent dynamics, ensemble composition, and sub-circuit, to the corresponding components as identified by CrEIMBO vs.
-other approaches.
-The sub-circuits were re-ordered for each method to maximize correlation with the ground-truth circuits using Scipy’s linear_sum_assignment [4].
+$\bm{F}^{\textrm{all}}=\arg\min_{\bm{F}^{\textrm{all}}}\|\bm{x}^{+}-\bm{F}^{\textrm{all}}(\bm{cx})\|_{2}^{2}+\lambda_{F}\|\text{vec}(\bm{F}^{\textrm{all}})\|_{1}+\lambda_{\rho}\sum_{{k_{1}},{k_{2}},({k_{1}}\neq{k_{2}})}\rho({\bm{f}}_{k_{1}},{\bm{f}}_{k_{2}})$ (3)
+where $\bm{F}^{\textrm{all}}\in\mathbb{R}^{p\times pK}$ is an horizontal concatenation of all $\bm{f}$s, and $\bm{x}^{+}\in\mathbb{R}^{p\times\sum_{d\in\textrm{batch}\mathrm{T}^{d}}}$ is the horizontal concatenation of all $\{\{\bm{x}_{t+1}^{d}\}_{t=1}^{\mathrm{T}^{d}}\}_{d\in\text{batch}}$, where “batch” means that this operation is in practice taken on a random subset of sessions for computational complexity considerations.
+Above, $\bm{cx}_{t}^{\textrm{batch}}\in\mathbb{R}^{Kp\times 1}=[(\bm{c}^{d}_{t}\otimes[1]_{1\times p})\circ([1]_{1\times K}\otimes\bm{x}^{d}_{t})^{T}]^{T}$, such that $\bm{cx}\in\mathbb{R}^{Kp\times\sum_{d\in\text{batch}}\mathrm{T}^{d}}$ is the horizontal concatenation of all $\{\{(\bm{cx})_{t}^{d}\}_{t=1}^{\mathrm{T}^{d}}\}_{d\in\text{batch}}$.
+The operator $\text{vec}(\cdot)$ flattens a matrix to a vector and $\lambda_{F}$ is $\ell_{1}$ sparsity-promoting weight on $\bm{F}^{\textrm{all}}$.
+$\rho(\cdot)$ represents correlation, such that $\sum_{{k_{1}},{k_{2}}}\rho(\bm{f}_{k_{1}},\bm{f}_{k_{2}})$ is a de-correlation term with weight $\lambda_{\rho}$ used to ensure that distinct $\bm{f}$s are not too similar.
+See Algorithm 1 for a method summary and Section E for complexity analysis.
+Figure 2: A-E: Testing CREIMBO over 312 repeats with varying random initializations and random seeds (Sec.
+B) reveals high correlations between ground truth and fit components for all unknowns, $\bm{Y}$, $\bm{A}$, $\bm{x}$, $\bm{F}$, $\bm{c}$ (subplots A-E respectively).
+F-G:Ablation Experiment: Comparison of the ensemble trajectory ($\bm{x}^{d}$, subplot F) and the ensemble compositions ($\bm{A}$, subplot G) for the $1$-st session as identified by CREIMBO vs.
+various ablations (details in Sec.
+D, see Figure 7 for more sessions).
+H-J: Comparing CREIMBO to ablation variants in terms of $-\log(\text{MSE})$ between ground truth and identified dynamics, ensembles, and sub-circuits.
+“Single Session” refers to training CREIMBO on sessions individually.
+Details in Sec.
+D.
+K: Baseline Comparisons: the latent dynamics identified by the baselines (SDLS, rSLDS (Linderman et al., 2016), multi-population rSLDS (Glaser et al., 2020), details in Sec.
+D) for the first two sessions (merged sessions baselines provide a single trajectory).
+L-O: Comparison of the latent dynamics and identified ensemble matrices compared to the ground truth across the baselines.
 Figure 3: CREIMBO identifies cross-regional neural sub-circuits underlying multi-session human brain recordings.
@@
 D: The standard deviation (std) of the coefficients over time across the two exemplary sessions.
-E: Examplary identified sparse ensemble matrices ($\{\bm{A}^{d}\}_{d=1}^{D}$) for the observations from A.
+E: Exemplary identified sparse ensemble matrices ($\{\bm{A}^{d}\}_{d=1}^{D}$) for the observations from A.
 F: Relative std of the coefficients across four subjects reveals usage of similar sub-circuits.
@@
 Black arrows near $\bm{f}_{8}$ and $\bm{f}_{11}$ highlight within-region interactions.
-I: Pairwise correlations between within-session sub-circuit coefficients (left, middle) and average within-session correlations across all sessions with std marked by the size of a “+” marker.
-J: Average and std of connection values in $f$s.
-K: Number of non-zero connections per sub-circuit $\bm{f}_{k}$.
+I: Pairwise correlations between within-session sub-circuit coefficients (left, middle) and average within-session correlations across all sessions.
+STD marked by “+” marker size.
+J: Mean and STD of connections $f$s.
+K: Number of non-zero connections per $\bm{f}_{k}$.
 4 Experiments
-4.1 CrEIMBO recovers Ground Truth Components in Synthetic Data
-We first assess CrEIMBO’s ability to recover ground-truth components from synthetic data.
+CREIMBO recovers Ground Truth Components in Synthetic Data: We first assess CREIMBO’s ability to recover ground-truth components from synthetic data.
 We generated $K=3$ sub-circuits represented by a set of 3 rotational matrices (Fig.
-4B).
+5B).
 These sub-circuits capture interactions among $p=3$ ensembles from $J=3$ regions ($p_{j}=1$ for all $j=1,\dots,J$).
@@
 We finally generated the data by simulating 5 sessions with different neuron identities and varying numbers of neurons per region(Fig.
-4A and D).
+5A and D).
 CREIMBO was able to recover the ground truth components with high accuracy, as measured by correlation with the ground truth, across $312$ random noisy initialization repeats with random parameters from a range described in 1.
@@
 2E).
-Comparing CrEIMBO’s results to components identified by other methods (baselines detailed in F), we found that approaches that either separate the ensembles and dynamics identifications steps, analyze each session separately, or do not consider the localized structure of the ensembles, fail to recover the ground-truth components, i.e., the resulting fit model parameters are less correlated with the ground-truth observations (Fig.
-2F-J, 5).
+Comparing CREIMBO’s results to components identified under diverse ablations (details in D), we found that approaches that either separate the ensembles and dynamics identifications steps, analyze each session separately, or do not consider the localized structure of the ensembles, fail to recover the ground-truth components, i.e., the resulting fit model parameters are less correlated with the ground-truth observations (Fig.
+2F-J, 7).
+We further evaluated CREIMBO’s ability to recover hidden components compared to other methods, (Fig.
+2 K-O, Fig.8.
+baselines details in Sec.
+D).
+While the latter shows improved performance, CREIMBO more accurately recovers the ground truth hidden dynamics and ensembles.
+Importantly, these models are not designed to identify multiple co-occurring processes or sparse ensembles, in contrast to CREIMBO, and they are not suitable for multi-sessions with varying neural identities, which limits the comparison from the outset.
 Next, we tested CREIMBO on a richer synthetic example with $D=15$ sessions.
 Each session had a varying number of neurons (a random integer between $14$ and $19$ per region from a uniform distribution) and distinct sub-circuit coefficients (see Fig.
-8 for the generated ground truth components and 6 for regional distributions and masking).
+11 for the generated ground truth components and 9 for regional distributions and masking).
 The sub-circuits were set as rotational matrices, similar to the previous experiment.
 Over $204$ random initializations with different parameters (see Sec.
-D) we again found that CrEIMBO robustly identifies the ground truth latent dynamics, ensemble compositions, and sub-circuits.
+B) we again found that CREIMBO robustly identifies the ground truth latent dynamics, ensemble compositions, and sub-circuits.
 Moreover, CREIMBO accurately reconstructed the observations with high correlation to the ground truth across all initializations (Fig.
-8E) with low relative error (Fig.
-8F).
+11E) with low relative error (Fig.
+11F).
 We further tested CREIMBO on more advanced data that included $D=40$ sessions with a maximum of $4$ regions and $p_{j}=3$ ensembles per region (Fig.
-9).
+12).
 We found that CREIMBO manages to recover the components with high correlation with the ground-truth (Fig 13).
-4.2 CrEIMBO Discovers Multi-Regional Dynamics in Human Electrophysiology Data
-We then tested CrEIMBO on human neural recordings from a high-density electrode array provided by [12].
-The data consists of neural activity from $p=10$ brain areas in the human medial temporal and medial frontal lobes, recorded while subjects performed a screening task (detailed in [12]).
+CREIMBO Discovers Multi-Regional Dynamics in Electrophysiology Data: We then tested CREIMBO on human neural recordings from a high-density electrode array provided by Kyzar et al.
+(2024).
+The data consists of neural activity from overall $p=10$ brain areas with limited cross-session regional overlap (Fig.
+14, 15), recorded while subjects performed a screening task (details in Kyzar et al.
+(2024)).
 The data encompasses $21$ subjects across $D=41$ non-simultaneous sessions.
-For each subject (except Subject 19), the data offers two recording sessions: 1) a “Screening” session to identify natural images that elicit strong neural responses for that subject and 2) a “Sternberg test” session to assess working memory using the images identified in the screening session.
-For Subject 19, the data includes only the Sternberg session.
-The data encompasses multiple regions with limited cross-session regional overlap, a few missing samples per session, and variability in the distribution and count of neurons over areas, with some areas having no neurons in some trials or subjects (Fig.
-11, 12).
-We first converted the spike-sorted units to estimated firing rate matrices by convolving the spike trains with a 30 ms Gaussian kernel (Fig.
-7).
-Applying CrEIMBO to the firing rate data with a maximum of $K=12$ sub-circuits and $p_{j}=7$ ensembles (Fig.
-13) per region, we found multi-regional motifs with variability across sessions.
-We then tested CrEIMBO on the first $18$ seconds from whole sessions together with a maximum of $K=12$ sparse sub-circuits and as most $p_{j}=7$ ensembles per region (full parameter list available in Tab.
+For each subject (except Subject 19), the data offers 1) a “Screening” session, and 2) a “Sternberg test” session (details in Kyzar et al.
+(2024)).
+We first converted the spike-sorted units to firing rate matrices by convolving the spike trains with a $30$ms Gaussian kernel (Fig.
+10), and then tested CREIMBO on all sessions together with a maximum of $K=12$ sub-circuits and as most $p_{j}=7$ ensembles per region (full parameter list in Tab.
 2).
@@
 Interestingly, the ensemble compositions identified by CREIMBO (Fig.
-3E) presented sparse patterns with one dense ensemble for most regions that we hypothesize captures the ‘general’ mean field activity of that region (the right-most column of each diagonal block in Fig.
+3E) present sparse patterns with one dense ensemble for most regions that we hypothesize capture the “background” mean field activity of that region (right-most column of diagonal blocks in Fig.
 3E).
 The sparse ensembles we hypothesize capture the more specialized, nuanced functionality.
-The ensemble dynamics ($\bm{X}^{d}$, Fig.
-3G) reveal that some ensembles are constantly active, implying that they might play an important role in the observed neural activity.
-The identified set of global sub-circuits ($\{\bm{f}_{k}\}_{k=1}^{K}$) exhibits distinct localized motifs (Fig.
+The ensembles activity ($\bm{X}^{d}$, Fig.
+3G) reveal some constantly-active ensembles, implying these neuron groups are important for neural processing.
+The universal sub-circuits ($\{\bm{f}_{k}\}_{k=1}^{K}$) exhibit distinct localized motifs (Fig.
 3H), i.e., most sub-circuits present clear trends of either sourcing or targeting the same area.
 This is in contrast to other approaches (Sec.
-F), that yield overlapping sub-circuits (Fig.
-19) that emphasize only limited circuitry (Fig.
-20).
-Interestingly, sub-circuits 10 and 12 ended up almost empty (mainly zero-ish values, Fig.
-3H, 16).
-This highlights CrEIMBO’s ability to automatically nullify extra or redundant sub-circuits—i.e., a form of model selection—via the sparsity regularization.
+D), that yield overlapping uninterpretable sub-circuits (Fig.
+22) that emphasize only limited circuitry (Fig.
+24).
+Interestingly, sub-circuits $10$, $12$ ended up almost empty (mainly zero-ish values, Fig.
+3H, 19).
+This highlights CREIMBO’s ability to automatically nullify redundant sub-circuits—i.e., a form of model selection—via the sparsity regularization.
 The number of active connections and average connection strength (Fig.
-7J, K) present modest variability between sub-circuits.
-This underscores the model’s ability to identify distinct sub-circuits with different structural behaviors, capturing various interaction resolutions.
-Specifically, it can find sub-circuits of varying strengths and is not limited to only k-sparse sub-circuits.
-From a neuroscience perspective, this suggests that underlying sub-circuits can vary in density and the number of participating areas and ensembles.
-Interestingly, most identified interactions occur between distinct areas (i.e., inter-region interactions) with either multiple ensembles from the same area affecting together ensembles from a different area (e.g., $\bm{f}_{3}$), or multiple ensembles from diverse areas converging to the same target area (e.g., $\bm{f}_{5}$).
-However, these source or target groups of ensembles vary across sub-circuits, hence the full repertoire of activity cannot be captured by a single.
-In addition to these between-area interactions, we also observed, although to a lesser degree, within-area interactions (e.g., as in $\bm{f}_{8}$ and $\bm{f}_{11}$, as marked by black arrows).
-These dynamics emphasize that within-area amplification or regulation dynamics do exist in addition to global cross-region computations, however seem to be less salient than cross-area communication.
-When exploring the activation patterns of these sub-circuits (Fig.
+10J, K) show modest variability between sub-circuits, highlighting the model’s ability to identify distinct sub-circuits with varying interaction resolutions, including those of varying strengths rather than being limited to just k-sparse sub-circuits.
+From a neuroscience perspective, this suggests that underlying neural sub-circuits can vary in density and the number of participating areas and ensembles.
+Interestingly, most identified interactions occur between distinct areas (i.e., inter-regional interactions) with either multiple ensembles from the same area affecting together ensembles from a different area (e.g., $\bm{f}_{3}$), or multiple ensembles from diverse areas converging to the same target area (e.g., $\bm{f}_{5}$).
+However, these source or target groups of ensembles vary across sub-circuits, hence the full repertoire of activity cannot be captured by a single one, highlighting the importance of CREIMBO’s capacity to disentangle the combined activity of multiple circuits.
+We also observe, but to a lesser degree, within-area interactions (e.g., black arrows in $\bm{f}_{8}$, $\bm{f}_{11}$).
+These dynamics emphasize that within-area amplification or regulation do exist in addition to global cross-region computations.
+When exploring the activation patterns of these identified sub-circuits (Fig.
 3D, Fig.
-17), we reveal variability between subjects and sessions.
-However, the usage of overall similar sub-circuits, with some sub-circuits not being used at all in any of the exemplary sessions (Fig.
-3C, F, e.g., $\bm{f}_{12}$ not being used), while others are constantly used with high variability over time (e.g., $\bm{f}_{8}$ light-green).
-Interestingly, when exploring the within-session pairwise correlations between sub-circuits coefficients (Fig.
-7I), it seems that most coefficients present very low ($<0.1$) correlations, indicating that the activity of the corresponding circuits is not similar and they might capture different cognitive processes, while other pairs of sub-circuits may show similarities in activity.
-This might indicate that for a specific subject or session, they may be related.
-For instance, a specific image may evoke some emotional response in one subject but not in another (e.g., $\bm{c}_{9}$ $\bm{c}_{11}$ in the second session of subject 10, however not in the first session of that subject nor in the average correlations across all sessions).
-5 Discussion.
-In this paper we introduced CrEIMBO (Cross Ensemble Interactions in Multi-view Brain Observations), a novel approach for uncovering multi-regional dynamics in neural data collected across multiple sessions.
-CrEIMBO addresses the challenge of integrating non-simultaneous neural recordings by leveraging joint inference and sparsity regularization to capture the underlying neural sub-circuits governing brain activity.
-CrEIMBO offers several advantages over other models of whole-brain multi-session data.
-Chiefly, it identifies a cross-session shared latent space where non-stationary ensemble interactions are governed by a set of global sub-circuits.
-By structuring the time-varying dynamics in terms of a basis set of sub-circuits, CrEIMBO allows for the exploration of active sub-circuits across sessions and aiding in the identification of subject-to-subject and cross-session variability in neural activity.
-Additionally, constraints on the loading matrices $\bm{A}$ enable the identification of area-specific ensembles, facilitating the understanding of within-area compared to between-area interactions.
-Moreover, the simultaneous fitting of the ensembles and dynamics automatically aligns cross-session ensembles in terms of functionality, which is critical for studying cross-session variability.
-We demonstrated the efficacy of CrEIMBO through both synthetic experiments and neural recordings.
-On synthetic data, CrEIMBO consistently recovered ground truth components with high accuracy.
-In our neural recordings analysis, CrEIMBO found cross-regional motifs that span interactions across sessions and subjects.
-Limitations and future work: While CrEIMBO leverages information from multiple brain sessions, the most modern Electrophysiology tools are still far from observing every neuron or region simultaneously.
-Consequently, and as with all dynamical systems models, obscured interactions with missing (i.e., unrecorded) neurons within the brain may limit the precision of the identified interactions.
-Additionally, CrEIMBO’s current implementation does not include an offset term.
-Adding this would increase complexity and require thorough exploration; hence, we will address it in future work.
-We limit CrEIMBO here to linear projections from the ensemble-space to per-session observation space.
-Extending this aspect of the model to more flexible, non-linear projections while maintaining interpretability is challenging and a future goal.
-CrEIMBO’s effectiveness in aligning ensembles across sessions relies on the assumption that these sessions share commonalities in their active sub-circuits.
-If sessions operate with entirely different sub-circuits, cross-session ensemble alignment is not guaranteed.
-Another limitation is that we assume Gaussian observation statistics, i.e., spiking data must be converted to rates to use CrEIMBO.
-Future research could incorporate non-Gaussian likelihoods, such as Poisson for spike counts.
-Finally, CrEIMBO rests on the mathematical foundation of dictionary learning, which can be computationally intensive at higher dimensions.
-To enhance computational efficiency, we plan to implement parallel processing, data batching, and GPU processing in future iterations.
-We also look to leverage CrEIMBO to identify pathological or cognitive states during which specific sub-circuits manifest exclusively (e.g., neural circuitry changes during epileptic attacks).
-Another extension for the current method includes extending the global cross-session circuits set to enable small, controlled structural adaptations of the sub-circuits between sessions, e.g., through an $\ell_{2}$ regularization on the distance between cross-session corresponding sub-circuits $\|\bm{f}_{k}^{d}-\bm{f}_{k}^{d^{\prime}}\|_{F}^{2}$.
-Acknowledgments
-N.M. was supported by the Kavli Discovery Award of the Kavli Foundation.
+20), CREIMBO reveals cross-subject and cross-session variability.
+Notably, some sub-circuits are consistently used with high variability over time (e.g., $\bm{f}_{8}$ light-green), while others, such as $\bm{f}_{12}$, are not utilized at all in any exemplary session (Fig.
+3C, F, e.g., $\bm{f}_{12}$).
+Moreover, when exploring sub-circuits with high activity variability across sessions (Fig.
+3D,F), we observe that some are used within-subjects but across both tasks (e.g., $\bm{f}_{2}$ in subject 14, Fig.
+3D), while others are shared across both subjects and tasks (e.g., $\bm{f}_{11}$, used by subjects 10 and 14 in both task phases).
+The analysis of within-session pairwise correlations between sub-circuit coefficients (Fig.
+10I) reveals low correlations ($<$0.1) for most pairs.
+This indicates that the activities of distinct sub-circuits differ and reflect different cognitive processes.
+Since ground-truth components are unavailable in real data, direct validation of such components is impossible.
+Hence, we sought a proxy to evaluate CREIMBO’s performance on real-data by testing its consistency under increasing observation noise levels.
+Particularly, we re-ran CREIMBO on the human recordings after adding increasing variance i.i.d Gaussian noise (with $K=8,p_{j}=6$ Fig.23).
+CREIMBO remains robust to increasing noise, and experiences a phase transition at a specific noise level ($\frac{\sigma_{\text{noise}}}{\sigma_{\text{data}}}\sim 0.2$,Fig.23 A,B, E, F,H,I), which align with the dictionary-learning literature (Studer & Baraniuk, 2012).
+In this noise experiment, CREIMBO further identifies that within-area ensemble composition correlations (Fig.23C,D) remain stable until a noise level of around $\sigma_{\text{noise}}\approx 10$, after which these correlations weaken as noise increases, which may imply that high noise in data can obscure meaningful relationships between ensembles.
+Additionally, it identified that the dynamics’ coefficients remain robust until $\sigma_{\text{noise}}\approx 1.8$, while exhibiting an increasing internal frequency as the noise increases, potentially to account for noise that does not follow dynamical rules (Fig.
+23G,  25).
+This suggests that frequency of fast transitions in CREIMBO’s ${\bm{c}_{k,t}}$ could indicate noise in the system that does not adhere to the underlying dynamics.
+CREIMBO Discovers Regional Interactions Predictive of Task Variables: We tested CREIMBO’s ability to infer task-related variables from mice whole-brain Neuropixels multi-session data during a memory-guided movement (data from Chen et al.
+(2024), Fig.
+2627).
+CREIMBO identifies intra- and inter-area brain interactions via the sub-circuits (Sec.
+F, Fig.
+29C,D), including cross-regional flows into or from key areas responsible for memory (from $\bm{f}_{5}$ and to $\bm{f}_{8}$ the hippocampus), planning (from frontal cortex, e.g.
+$\bm{f}_{3}$), and movement (from primary motor cortex, $\bm{f}_{4}$), which are needed for the task.
+Within-area sub-circuits (Fig.
+29D) further showed within-area ensemble interactions (e.g., the basal ganglia in $\bm{f}_{1}$ and the secondary motor cortex in $\bm{f}_{2},\bm{f}_{7}$, along with self-activation/inhibition of ensembles in other sub-circuits (e.g.
+$\bm{f}_{4}$).
+Moreover, CREIMBO’s dynamic coefficients (${\bm{c}_{kt}}$) captured task-related patterns across trials and sessions (Fig.
+4A,B).
+When training a regularized logistic-regression model, we found that we could predict from CREIMBO’s circuits activations ($\bm{c}_{kt}$) various task variables including outcome, early lick, and lick side far above chance levels (Fig.
+4C-J, p-values < $1\times 10^{-10}$).
+Furthermore, CREIMBO’s sub-circuit activations across different task periods highlight how specific multi-regional interactions capture different aspects of the task.
+For instance, towards the trial end ($t_{3}$ window), which includes the lick movement, $c_{7}$ shows increased importance (Fig.
+4E).
+Notably, $c_{7}$ corresponds to the activity of the $\bm{f}_{7}$ sub-circuit that captures flows to the secondary motor cortex (Fig.
+29C).
+Another example is $c_{8}$ (activity of sub-circuit $\bm{f}_{8}$ that includes flows into hippocampus 29C), which shows increased feature importance in the first time window $t_{0}$.
+This aligns with the potential involvement of memory-related processes required with the stimulus appearance.
+Thus CREIMBO demonstrates the ability to capture interpretable neural dynamics and predict task variables that reveal complex regional interactions.
+(see Sec.
+F for more detailed results).
+Figure 4: Results of task-variable prediction using CREIMBO’s dynamic coefficients as input.
+Prediction was done based on coefficients from all used trials and sessions (19 sessions, 40-60 trials each, see Sec.
+F).
+A: Dynamic coefficients from example traces, shade marks within trial, bordered by starting and ending point.
+B: Dynamic coefficients for correct (“hit”) vs.
+wrong (“miss”) trials.
+C: Accuracy score of predicting task variables based on coefficients vs.
+chance.
+D, F,G,H,I: Confusion matrices of predicting varying task variables based on dynamic coefficients.
+E: importance of the different coefficients and time points for prediction.
+5 Discussion, Limitations, and Future Work
+Here, we introduced CREIMBO—a novel approach for uncovering multi-regional dynamics in neural data collected across multiple sessions.
+CREIMBO addresses the challenge of integrating non-simultaneous neural recordings by joint dynamical inference and sparsity regularization to capture the underlying neural sub-circuits governing brain activity.
+We further demonstrated the efficacy of CREIMBO through multiple synthetic and neural data, and found that CREIMBO recovered ground truth components and is robust to noise in identifying cross-regional motifs that span cross-session interactions.
+CREIMBO offers several advantages over existing methods.
+Chiefly, it identifies a cross-session shared latent space where non-stationary ensemble interactions are governed by a time-varying decomposition of universal basis dynamics.
+By structuring these dynamics in terms of global sub-circuits, CREIMBO allows the discovery of sub-circuits meaningful for various cognitive processes, enabling the identification of variability in neural activity across subjects and sessions.
+Limitations & Future Directions: An important feature of CREIMBO is its ability to unify sessions with different neuron subsets through the universal dictionary of dynamical interactions prior that aligns ensembles in terms of functionality.
+This can enable the inference of their activity even if they are not observed in some sessions.
+This ability, however, depends on the overlap between the used basis interactions across sessions, the distinctiveness of different $\bm{f}_{k}$s, and the premise that the basis interactions capturing interactions with ensembles missing from a session include dynamic values for ensembles observed in that same session (Sec.
+Assumptions).
+Another limitation is CREIMBO’s reliance on linear projections from ensemble to session observation space, restricting flexibility.
+Extending to non-linear projections offers potential for development, though it introduces computational and interpretability challenges.
+Finally, CREIMBO uses dictionary learning, which is computationally demanding, and future iterations will include parallel processing.
+Extending CREIMBO to additional data types and applications, including non-neural data (e.g., immune-cell counts), or for identifying pathologies, is an exciting future step.
+Moreover, integrating CREIMBO with mDLAG (Gokcen et al., 2024), with the latter identifying optimal communication delays and the former leveraging these to further identify the underlying set of co-active LDSs, presents an exciting future work.
 Appendix A Notations
@@
 $\tau=\mathrm{T}_{m}^{d}$ Duration of the trial or session
-Appendix B Differences from dLDS
-While our proposed model draws inspiration from dLDS [16], it significantly diverges in its approach, particularly in ensemble identification and the details thereof.
-Below we list the key differences:
-  1. 1.
-    Incorporation of sparsity constraints on the operators $f$.
-  2. 2.
-    Extension to multi-sessions through shared $f$s and distinct projections ($D$) instead of coefficients ($c$).
-  3. 3.
-    Imposition of regional structure on $D$.
-  4. 4.
-    Integration of graph-driven sparsity on $D$.
-  5. 5.
-    Modifications to the inference procedure inspired by linocs.
-  6. 6.
-    Ability to capture inter-subject variability.
-  7. 7.
-    Adaptation to handle variations in trial duration, fluctuations in neuron count, and changes in distributions between sessions.
-Appendix C Differences from SiBBlInGS
-Although our proposed model draws inspiration from SiBBlInGS [17] for ensemble identification, it introduces several fundamental differences:
-  1. 1.
-    SiBBlInGS does not consider interactions between ensembles, whereas our model incorporates ensemble interactions.
-  2. 2.
-    SiBBlInGS does not address differences in the number of neurons between subjects, whereas our model accounts for this variability.
-  3. 3.
-    SiBBlInGS lacks consideration for region-specific structure, focusing primarily on a single block of $D$, whereas our model imposes region-specific structure on $D$.
-  4. 4.
-    SiBBlInGS does not account for the evolution of the ensemble activity space, whereas our model incorporates this aspect.
-Appendix D Parameters Range for Synthetic Data Experiment
-The full set of parameters for the synthetic experiments are available in Table 1.
+Appendix B Synthetic Data Experiment Details
+We simulated the synthetic data to represent the non-stationary dynamics of brain activity including multi-region interactions.
+Current assumptions about brain activity suggest Zeki (2015); Nelson & Bower (1990) it functions through distributed, parallel processing across multiple regions, potentially involving co-active circuits that can process a range of variables (e.g., task demands, feedback signals, and sensory information), simultaneously.
+This biological reality thus motivated our data generation approach, to reflects the multi-process nature of neural activity, where different processes can interact and contribute to brain dynamics.
+We simulated $D=5$ sessions, each sharing a common set of basis rotational dynamics, denoted as $\{\bm{f}_{k}\}_{k=1}^{K=3}$, with $p_{j}=3$ regions of interest.
+Different regions were associated with different ensembles, and the dynamics for each session were modeled using sparse decomposition of the basis dynamics.
+For each session, the dynamics’ coefficients, $\bm{c}_{t}^{k}$, were generated over $T=500$ time points.
+These coefficients were generated by initially giving them binary values (active vs.
+inactive states), with random switches occurring every $x$ time points, where $x$ was uniformly drawn between 10 and 30.
+The sparsity constraint was maintained at each time point, ensuring that only one dynamic was active at any time ($\|\bm{c}_{t}^{k}\|_{0}=1\forall t$).
+Then, to introduce smooth transitions between switching states, a Gaussian convolution with a time window of 6 points (standard deviation = 0.7) was applied to the coefficients, to promote a smoother / more gradual changes between active states.
+Different sessions maintained the same sequence of active dynamics but allowed for varying durations of each state, introducing temporal variability across sessions.
+Each session $d$ was augmented with a unique projection matrix $\bm{A}^{d}$ that captures the mapping between the latent space and the observations (neural) space in the respective regions.
+These projection matrices $\bm{A}^{k}$ consisted of $P=3$ columns, each corresponding to one of the three regions.
+The number of neurons in each region was randomly chosen between 4 and 9, with the count varying independently across regions and sessions.
+This variability ensured that different sessions were not identical in their neural configurations.
+Latent dynamics for each session were generated using the recursive equation $\bm{x}_{t}=\sum_{k=1}^{K}\bm{c}_{kt}\bm{f}_{k}\bm{x}_{t-1}$, starting from an initial state $\bm{x}_{0}=[1,-1,1]^{\top}$.
+CREIMBO was trained based on the observations only, which were defined for each session as $\bm{y}^{d}=\bm{A}^{d}\bm{x}^{d}+\epsilon$ where $\epsilon$ is an addition of i.i.d Gaussian noise with standard-deviation of $\sigma_{\text{noise}}=0.1$
+The full set of parameters for the synthetic experiment are available in Table 1.
 Table 1: Parameter values for synthetic experiments
@@
 lasso_solver ’spgl1’ Pylops solver used for $\ell_{1}$ regularization.
-Appendix E Parameters for human multi-regional experiment
+Appendix C Parameters for human multi-regional experiment
 The full set of parameters for the human data experiment are available in Table 2.
@@
 l1_D 2.74 $\ell_{1}$ regularization for $\bm{A}$.
-Appendix F Information about the baselines:
+Appendix D Information about the baselines:
+We distinguish between two types of baselines: (1) ablation experiments and (2) comparison to the performance of entirely different methods.
+It is important to note that the latter cannot fully capture the unique capabilities of CREIMBO, particularly in identifying basis ensemble interactions shared across sessions through sparse decomposition.
+Therefore, we only compare certain aspects that can be extracted from other methods (e.g., multi-regional SLDS, which requires the same number of neurons per session).
+However, these comparisons do not imply that other models outperform or underperform CREIMBO; they simply offer different insights, with varying strengths and limitations depending on the application.
+Importantly, since the sub-circuits $\{\bm{f}_{k}\}$ is invariant to its ordering, we used SciPy’s implementation of the ‘linear_sum_assignment‘ problem Crouse (2016) to re-order the identified sub-circuits before comparing them with the ground truth, matching the $\bm{f}$s values by minimizing the $\ell_{2}$ error.
+Ablation Experiments:
   1. 1.
@@
   3. 3.
-    All Regions (sparse)”: Similar to the previous one but without applying sparsity on the ensemble matrix.
+    All Regions (non-sparse)”: Similar to the previous one but without applying sparsity on the ensemble matrix.
   4. 4.
@@
     Single Session #”: Running CREIMBO on a single session (view) of the data rather than leveraging cross-session information.
-    Since $\{\bm{f}_{k}\}$ is invariant to its ordering, we used SciPy’s implementation of the ‘linear_sum_assignment‘ problem [4] to re-order the identified sub-circuits before comparing them with the ground truth, matching the $f$ values by minimizing the $\ell_{2}$ error.
-Appendix G Code and Data Sharing
-The neural data used in this study [12] is publicly available with License: “spdx:CC-BY-4.0”.
-We used version 0.240123.1806 of the data which can be accessed through the DANDI Archive at https://dandiarchive.org/dandiset/000469/0.240123.1806 (Access Information: dandi:OpenAccess).
-The code will be shared via GitHub upon publication.
-Appendix H Technical Details
-The whole code is written in Python.
-We utilized the following Python packages along with their respective versions: scipy (1.8.0), matplotlib (3.8.2), numpy (1.23.5), pandas (1.5.0), seaborn (0.11.2), scikit-learn (1.0.2), pylops (1.18.2), and statsmodels (0.13.2), all within a Python 3.10.4 environment.
-For the $\ell_{1}$ regularization, we used the “SPGL1” solver of pylops [27].
-During the writing process, we used ChatGPT very sparingly to improve wording and fix grammar mistakes at the word or sentence level.
-The synthetic experiment can run locally on a desktop (simple ASUS machine).
-However, the robustness tests with hundreds of random repeats and the human data experiments used the entry node with CPUs only of NERSC’s “Perlmutter” supercomputer.
-Running time for each synthetic instance ranges from 0.8 to 2 minutes (with the capability to run in parallel).
-For the human data, processing all 41 sessions with 1000 time points per session (and varying neuron counts as described in the text and figures) takes between 17 to 25 minutes, depending on parameters and result quality.
-Appendix I Societal Impact and Potential Harmful Consequences
-With respect to Societal Impact, the development, application, and future work of CrEIMBO can advance our understanding of brain function.
-Future work can potentially lead to improved diagnostics of neurological disorders.
-We believe our method does not present any significant harmful consequences.
-We explicitly discussed the method’s limitations, ensured proper use of the data, and adhered to the NeurIPS code of ethics.
+Other methods:
+  1. 1.
+    SLDS”: We used the SSM Python package described by Linderman et al.
+at Linderman et al.
+(2020).
+For the (non-recurrent) SLDS option, we used the “gaussian_orthog“ emission parameter, the “bbvi” as the fitting method, “variational_posterior” was set to “mf”, and the number of iterations (“num_iters”) was set to 500.
+The rest of the fitting parameters were left as default.
+We were inspired by this notebook  https://github.com/lindermanlab/ssm/blob/master/notebooks/3-Switching-Linear-Dynamical-System.ipynb by  Linderman et al.
+(2020) for our comparison.
+This includes:
+    1) “SLDS (trials)” which captures the training of SLDS with “bbvi” fitting method from across all trials (by stacking trials information vertically),
+    2) “SLDS (per-trial)” which refers to training SLDS individually per-trial.
+  2. 2.
+    rSLDS”: Similarly to SLDS, for the recurrent version (rSLDS, (Linderman et al., 2016)), we used the same SSM package provided by Linderman et al.
+(2020).
+Here, we used the recurrent option for the transitions (i.e., the “transitions” parameter was set to“recurrent_only”), using “diagonal_gaussian” dynamics and “gaussian_orthog” emissions.
+This includes:
+    3) “rSLDS (trials)” which captures training rSLDS with the “bbvi” fitting method from across all trials (by stacking trials information vertically),
+    4) “rSLDS (per-trial)” which refers to training rSLDS individually per-trial.
+  3. 3.
+    “mp_rSLDS”: Multi-Regional rSLDS, as described in Glaser et al.
+(2020).
+To run our comparison, we were inspired by the colab-notebook for this “mp_rslds” method, provided by the SSM Python Package Linderman et al.
+(2020) (at https://colab.research.google.com/github/lindermanlab/ssm/blob/master/notebooks/Multi-Population-rSLDS.ipynb).
+This includes:
+    5) “mp_rSLDS-Gauss (trials)” which captures training “mp_rSLDS” from across all trials (by stacking trials information vertically) under Gaussian statistics.
+    6) “mp_rSLDS-Gauss (per-trial)” which refers to training “mp_rSLDS” individually per-trial under Gaussian statistics.
+    7) “mp_rSLDS-Poisson (trials)” which captures training “mp_rSLDS” with the “bbvi” fitting method from across all trials (by stacking trials information vertically) under Poisson statistics.
+    8) “mp_rSLDS-Poisson (per trial)” which refers to training “mp_rSLDS” individually per-trial under Poisson statistics.
+Assumptions
+The CREIMBO model is based on a core set of assumptions over the nature of the data.
+These assumptions draw on both known properties of neural processes and general well-known statistical models used widely in data science.
+We categorize these assumptions into two types:
+  1. 1.
+    Priors over the underlying neural processes.
+  2. 2.
+    Priors over Observational Constraints in order for CREIMBO to properly infer the underlying system by leveraging multi-session information.
+D.1 Priors over the underlying neural processes
+  1. 1.
+    The neural dynamics in each session $d$ are assumed to lie on a low-dimensional manifold, embedded in a $P<<N^{d}$ low-dimensional space that is defined by $P$ functional groups of neurons (referred to as “ensembles”).
+  2. 2.
+    We assume that these functional ensembles consist of neurons with co-activation patterns, where neurons can belong to more than one, but only a few (a sparse number) of ensembles, with varying degrees of membership.
+  3. 3.
+    We further propose that the interactions between these ensembles drive the evolution of the latent manifold over time and are key to encoding changes in conditions and behavior.
+  4. 4.
+    For CREIMBO to be effective, we assume these interactions arise from the joint synchronous activity of multiple co-occurring processes, captured by a limited-size set of “basis-ensemble-interactions”.
+The time-varying, sparse decomposition of these basis-ensemble-interactions, weighted by their time-local contributions in every time point, can adequately describe the manifold’s evolution over time.
+  5. 5.
+    We assume that different dynamics basis elements capture distinct processes or behaviors.
+Some processes are required to be globally related to the cognitive task, while others capture session- or subject-specific processes (see Sec.
+D.2).
+  6. 6.
+    Each of these ensemble interactions may capture between-area interactions, within-area interactions, or both.
+D.2 Priors over Observational Constraints
+To effectively learn a unified representation by leveraging information across sessions, we make the following assumptions on the statistics of the data.
+  1. 1.
+    Our ability to infer the underlying latent state (see Fig.
+1B) is contingent upon the overall observability of the dynamical system.
+In traditional linear systems, the observability matrix can guarantee that any state is visible in a finite time horizon output.
+Non-stationary systems, as in CREIMBO, do not have succinct guarantees, the closest of which come in the form of observability conditions on switched linear systems (Tanwani et al., 2012).
+    As the heart of the observability condition is that the dynamics “well mixes” the state such that any state element in the null space of the readout matrix will eventually be rotated into the its span, and thus visible.
+We thus assume that the spectral radius of each dynamical system (a measure of mixing in linear systems (Simchowitz et al., 2018)) in the basis is close to one.
+This assumption is loose, due to the lack of theoretical guarantees on general nonstationary systems, and further analysis should identify tighter, and less stringent, assumptions.
+  2. 2.
+    Another prior we want to consider is which $\bm{f}$s capture ensembles unobserved in certain sessions, allowing us to assert that CREIMBO can recover the activity of these unobserved ensembles during those sessions.
+To achieve this, we must ensure that in the sessions where these ensembles are unobserved, at least some $\bm{f}$s representing them also capture the activity of observed ensembles.
+For example, if certain missing ensembles are represented solely by $\bm{f}$s that do not capture any observed ensembles, we will not be able to infer the activity of these unobserved ensembles in the sessions where they are absent, because there are no observed ensembles linked through the dynamic prior.
+  3. 3.
+    To ensure uniqueness of the sparse decomposition of the dynamical systems model at each session and time point, we assume that the effective spark of the dynamics basis is large ($S^{*}\geq 2S)$.
+Essentially, there is no $\bm{f}_{k}$ that can be linearly composed of $2S^{*}$ other dynamical systems.
+This ensures that there cannot exist two (or more) equivalent decompositions $\sum_{k=1}^{K}\bm{f}_{k}c_{kt}$ and $\sum_{k=1}^{K}\bm{f}_{k}c_{kt}^{*}$ for distinct coefficients that both compose the same dynamical matrix.
+  4. 4.
+    Lastly, to ensure accurate cross-session alignment, we propose that each dynamical system projects uniquely onto the obseverved subset of ensembles in each session.
+This assumption prevents two dynamical systems (e.g., “ensemble 1 $\rightarrow$ ensemble 2 $\rightarrow$ ensemble 3” and “ensemble 1 $\rightarrow$ ensemble 4 $\rightarrow$ ensemble 3”) from being indistinguishable from the data available in that session (e.g., a session with only ensembles 1 and 3 recorded).
+Appendix E Computational Complexity
+CREIMBO’s learning process involves learning both the ensemble compositions per session ($\{\bm{A}^{d}\}_{d=1}^{D}$), the ensembles’ temporal activity ($\{\bm{x}_{t}^{d}\}_{d=1}^{D}$), and its underlying temporal evolution, which requires identifying the global (session-invariant) dynamic operators ($\{\bm{f}_{j}\}_{j=1}^{J}$) and their per-session temporal coefficients ($\{\bm{c}_{t}\}_{t=1}^{T}$).
+E.1 Complexity of ensemble-matrix update
+The loading matrix update relies on 4 main computational steps:
+1) Channel Graph Construction: This operation, performed once for all $N$ channels of every state $d=1\dots D$, generates a channel graph $\bm{H}^{d}\in\mathbb{R}^{N\times N}$ for each state $d\in[1,D]$ by concatenating within-state trials $1\dots M_{d}$ horizontally, resulting in a $N\times\sum_{m=1}^{M_{d}}T_{m}^{d}$ matrix.
+For simplicity, let $\widetilde{T}=\sum_{m=1}^{M_{d}}T_{m}^{d}$.
+The computational complexity of calculating the pairwise similarities of this concatenated matrix for all $D$ states is thus $\mathcal{O}\left(D\widetilde{T}^{2}N(N-1)\right)$.
+2) The k-threshold step: involves keeping only the $k$ largest values in each row while setting the other values to zero—the complexity will be $\mathcal{O}\left(\widetilde{T}\log k\right)$ per row for a total computational complexity of $\mathcal{O}\left(DN\widetilde{T}\log k\right)$ for $N$ rows and $D$ states.
+3) State Graph Construction: This is a one-time operation that involves calculating the pairwise similarities between each pair of states.
+For simplicity, if we assume the case of user-defined scalar labels, and as in this case there are $D$ states (and accordingly $D$ labels), the computation includes $D(D-1))/2$ pairwise distances for $\mathcal{O}\left(D^{2}\right)$.
+4) Ensemble Inference (Eq.
+equation 1): This iterative step involves per-channel re-weighted $\ell_{1}$ optimization.
+If the computational complexity of a weighted $\ell_{1}$ is denoted as $\mathcal{C}$, then the computational complexity of the re-Weighted $\ell_{1}$ Graph Filtering is $NL\mathcal{C}+LNk$, where $N$ is the number of channels, $L$ is the number of iterations for the RWLF procedure, and $k$ is the number of nearest neighbors in the graph.
+For the last term in Eq.
+equation 1, there are $p^{2}$ multiplicative operations involving the vector $\nu$ and the difference in ensembles, arising from the $\ell_{2}^{2}$ norm.
+Additionally, there is an additional multiplication step involving $\bm{P}_{dd^{\prime}}$.
+For each state $d$, this calculation repeats itself $D-1$ times (for all $d^{\prime}\neq d$).
+This process is carried out for every $d=1\dots D$.
+In total, these multiplicative operations sum up to $\left(p^{2}+1\right)D(D-1)$, resulting in a computational complexity of $\mathcal{O}\left(D^{2}p^{2}\right)$.
+E.2 Complexity of inferring the ensembles activity and interactions coefficients (Eq. equation 2)
+The inference of $\bm{c}_{t}^{d}$ for each $t=1\dots T$ amd each condition $d=1\dots D$, involves: $\widehat{\bm{c}}_{t}^{d}=\arg\min_{\bm{c}_{t},\bm{x}_{t}}\|\bm{y}_{t}^{d}-\bm{a}\bm{x}_{t}\|_{F}^{2}+\|\widetilde{{\bm{x}}_{t+1}}^{d}-\bm{FX}_{K}^{d}\bm{c}_{t}^{d}\|_{F}^{2}+\lambda_{c}\|\bm{c}_{t}\|_{1}$ where $\bm{FX}\in\mathbb{R}^{p\times K}$ is the horizontal concatenation of $\{\bm{f}_{k}\bm{x_{t}}\}_{k=1}^{K}$.
+This simplifies to a Lasso problem of the form
+$\min_{\theta}\frac{1}{2}\|\xi-\mathcal{M}\theta\|^{2}\quad\text{s.t.}\quad\|\theta\|_{1}\leq\tau,$
+with dimensions $\xi:(N+p)\times 1$, $\mathcal{M}:(N+p)\times(p+K)$, and $\theta:(p+K)\times 1$, where
+$\xi=\begin{bmatrix}y_{t}\\ 0_{p\times 1}\end{bmatrix}\in\mathbb{R}^{(N+p)\times 1},\quad\theta=\begin{bmatrix}x_{t}\\ c_{t}\end{bmatrix}\in\mathbb{R}^{(p+K)\times 1},$
+and
+$\mathcal{M}=\begin{bmatrix}A&0_{N\times K}\\ I_{p}&\bm{FX}\end{bmatrix}\in\mathbb{R}^{(N+p)\times(p+K)}.$
+This can be solved with SPGL1 with up to num_iters iterations, resulting in a computational complexity of
+$O(\text{num\_iters}\cdot(N+p)(p+K)).$
+per session and time point and overall (across all sessions and time points):
+$O(\text{num\_iters}\cdot(N+p)(p+K)\sum_{d=1}^{D}T^{d}).$
+E.3 Complexity of inferring the dynamics’ dictionary
+The inference of the basis-interactions dictionary $\{\bm{f}_{k}\}_{k=1}^{K}$ (via Eq. equation 3) involves solving
+${\widehat{\bm{F}}_{all}}=\arg\min_{{\bm{F}}_{all}}{\|{\bm{x}}^{+}-\bm{F}_{all}\bm{({CX})}\|_{F}^{2}},$
+where $\bm{x}^{+}\in\mathbb{R}^{p\times\sum_{d\in\textrm{batch}\mathrm{T}^{d}}}$ and $\bm{cx}\in\mathbb{R}^{Kp\times\sum_{d\in\text{batch}}\mathrm{T}^{d}}$.
+For simplicity, assuming that $\lambda_{\rho}=0$, The problem simplifies to a LASSO formulation, which is solved using SPGL1 for $F^{\text{all}}$, with an overall computational complexity per iteration of $O(\text{num\_iter}\cdot Kp^{2}\tau)$, where num_iter represents the number of iterations required for convergence.
+Appendix F Application to multi-Regional cross-Session Mice Neural Activity
+Figure 26: number of neurons per area across all files
+Figure 27: number of neurons per area as a motivation for area-focusing.
+A: Overall number of neurons per sub-area (colored by parent area) across all sessions.
+B: Distribution of areas with sub-area name across all sessions.
+C: Distribution of areas with parent-area name across all sessions.
+D: Number of neurons per parent-area (across the 10 selected areas) in the 19 used sessions.
+Figure 28: Example Session  Chen et al.
+(2024).
+A: Data firing rate over time separated by trials in vertical lines.
+the color of the vertical line captures the relevant brain area (example random session).
+B: Trials start and end times (in seconds) for examples session.
+C: Trial duration distribution for example random session.
+Figure 29: Identified components of the Mesoscale data (Chen et al., 2024) using 19 random sessions with 40-60 trials within each.
+A: Identified ensemble matrices for three random sessions.
+B: Correlations between circuits.
+C: Between-areas interactions (extracted from $\{\bm{f}_{k}\}_{k=1}^{K}$ via a block off-diagonal mask) D: Within-areas interactions (extracted from $\{\bm{f}_{k}\}_{k=1}^{K}$ via a block diagonal mask)
+Figure 30: The components identified by the SLDS (Fig.
+30) and rSDLS (Fig.
+31) baselines (Linderman et al., 2016; 2020), when we applied them to the mouse mesoscale data from all sessions together (A) or individual sessions (B).
+For each model, we showed both the sub-circuits and their active coefficients for the first few trials.
+Compared to CREIMBO, the identified sub-circuits are more distributed and dense, and the dynamic coefficients capture only sharp transitions, i.e., unable to represent multiple co-occurring processes, in contrast to CREIMBO.
+Checking individual sessions with SLDS (B) or rSLDS (C), resulted in per-session set of sub-circuits (Fs) that cannot enable cross-session sub-circuits comparison.
+All models ran with $K=10$ sub-circuits and $p_{j}=3$ ensembles per-region, to keep maximum consistency with our CREIMBO results from Figure 29.
+Figure 31: Same as Figure 30, but for the rSLDS baseline Linderman et al. (2016).
+Table 3: Parameter table for the CREIMBO on Mice Mesoscale.
+Parameter Value
+max_error 1e-09
+seed 0
+normalize Fs True
+sparse_f True
+sparsity_on_f_max 40
+increase_in_sparsity_f 1.4318632048860938
+norm_D_cols True
+D_graph_params {’with_kNN’: True, ’with_norm’: True, ’k’: 15}
+lambda_D 0.4392235894758376
+update_type_D spgl1
+latent_dyns_initialization random
+D_with_lasso True
+l1_D 2.7440675196366238
+params_D {’update_c_type’: ’spgl1’}
+num_regions 10
+reg_type_on_c spgl1
+lambda_x 1.0291322176748077
+latent_dim 30
+num ensembles per region 30
+number of sub-dynamics 10
+seed_f 0
+sigma_mix_f 0.1
+sparse_f_params {’axis’: ’1’, ’percent0’: 20}
+We further tested CREIMBO’s ability to infer meaningful task variables underlying whole-brain multi-session data from mice Chen et al.
+(2023; 2024) while performing a memory-guided movement task.
+The measurements were obtained using multi-electrode extracellular electrophysiology, with the overall dataset stored in DANDI Archive under the NWB format Chen et al.
+(2023).
+Particularly, the data include electrophysiology recordings of neural activity from multiple brain areas per session (Figs.
+26,  27), across overall $175$ sessions, with numerous trials in each.
+The task the mice performed during the recordings involved selecting one of two “lick ports” based on auditory cues, with the left or right port leading to a reward depending on whether they heard a high or low tone.
+Neural activity was recorded across several brain regions in each session, with some regions overlapping between sessions, though no two sessions had identical coverage (refer to Fig.
+1., A-B, in Chen et al.
+(2024)).
+We first loaded data from all sessions and selected areas for CREIMBO that featured the highest number of sessions including them.
+We identified the parent area of each sub-area labeled in the dataset and removed neurons with unlocated or NaN areas.
+Inspired by our exploratory analysis of parent-area distributions across sessions and neurons (Fig.27B,C), we selected the 10 parent areas with the highest number of neurons across sessions [^1] Used Parent Areas: Thalamus, Orbital Cortex, Secondary Motor Cortex, Basal Ganglia, Olfactory Cortex, Brain Stem, Somatosensory Cortex, Frontal Cortex, Hippocampus, Primary Motor Cortex.
+We then selected 20 random sessions.
+One of these 20 sessions contained only ‘NaN‘-located neurons, leaving 19 sessions for this analysis (Fig.
+27D) [^2] Sessions DANDI names: ‘78a58614-09e4-4e29-a492-8bbeaf7a8cf6‘, ‘ab103954-d99f-46b1-98ac-e6f91a1e9313‘, ‘32701cfa-8932-4d9b-9f4d-cc05cb22ae89‘, ‘b080c738-17c9-4e0d-aee4-6d5371518f69‘, ‘886c4302-846a-4ef5-996a-6f02d6a81a5f‘, ‘ba5b1ff3-753e-425c-baf0-e7cdd0c08093‘, ‘ea856208-4240-404a-9034-d729ad6f4cda‘, ‘ed759efa-dceb-4472-a2a3-4f7357d67665‘, ‘8a2ce9b2-2e98-4c37-8f2a-3b9c8b542086‘, ‘22791d80-26dc-4495-b4c0-651fe10e3298‘, ‘a007fac1-96e7-4028-ad84-31a1b80db089‘, ‘b39eb52a-0774-411b-b374-6eb08a73562e‘, ‘8dbc25ee-cc17-4504-a1b9-7d43643b9466‘, ‘981feb84-f209-4994-8838-02a0048feb87‘, ‘585dc9d4-be9c-4244-8fa7-e64309019afc‘, ‘aa74e4d7-a79b-4179-adf3-497fb1237edc‘, ‘ebfab58b-7c80-4f31-9b4e-2b33883bb14a‘, ‘8c724097-3876-48fc-94e0-832779e4f1be‘, ‘2e9cb8a1-457d-49ea-97fd-f7023434d231‘.
+We pre-processed the data using a 30 ms Gaussian kernel for firing rate estimation across all the sessions used (see  2).
+We then selected only the neurons from the 10 parent areas and limited the analysis to the first 60 trials.
+Some sessions exhibited no activity or disconnection in certain trials, meaning some sessions included fewer than 60 trials (but no fewer than 40).
+Next, we ran CREIMBO with 10 sub-circuits ($K=10$) and $p_{j}=3$ ensembles per area.
+The full list of parameters is shown in 3, and an example exploration of a random session is shown in (Fig.28,A-C show the firing rate, trial start-end times, and distribution of trial durations, respectively).
+We identified ensembles per region (exemplary in 29A) with sub-circuits that capture both between-area (Fig.29C) and within-area interactions (Fig.29D), showing diverse motifs with varying degrees of cross-circuit correlations (Fig.
+29B).
+Some cross-area interactions reveal effects originating from specific regions involved in the task.
+For example, $\bm{f}_{1}$ demonstrates hippocampal effects, which align with the task’s memory demands, while, e.g., $\bm{f}_{3}$ shows interactions involving the frontal cortex, including follow-up activations, corresponding to the task’s need for planning and execution.
+Other sub-circuits reveal inputs converging on specific areas.
+For instance, $\bm{f}_{8}$ shows motor cortex inputs, relevant for moving the tongue during the lick response, and $\bm{f}_{2}$ reflects inputs into the secondary motor cortex, which is involved in coordinating motor actions.
+When extracting only within-area interactions, we observe unique motifs as well.
+For example, $\bm{f}_{1}$ shows further internal processing between basal ganglia ensembles, which may reflect the internal coordination and integration of motor planning and execution signals.
+$\bm{f}_{7}$ and $\bm{f}_{2}$ show strong effects of one ensemble on another within the secondary motor cortex, which could represent the modulation and refinement of motor commands for more precise movements.
+$\bm{f}_{5}$ demonstrates brainstem effects, with two ensembles influencing two other ensembles, potentially supporting basic motor functions such as muscle activation and coordination.
+$\bm{f}_{4}$ shows multiple self-activations of ensembles, which may capture self-regulatory processes like feedback inhibition or facilitation that modulate motor output and prevent over-activation.
+We further examined CREIMBO’s dynamics coefficients (${\bm{c}_{t}}$) and identified task-related patterns with similar coefficient profiles across trials for the same task variable (Fig 4A for an example session, and B for hit vs.
+miss outcomes).
+Similar differences were observed across other task variables.
+To quantify the predictive power of the dynamic coefficients, we used them as input for training a simple one-vs-rest logistic regression classifier.
+Each trial was split into four equal-duration time windows, and the 10 dynamic coefficients were averaged within each window, resulting in 40 features (4 windows × 10 coefficients).
+These features were then used to predict task variables, including outcome (hit/miss/ignore), early lick (whether a lick was early), lick side instructed, and lick side performed in practice.
+Additionally, we tested the model to predict 2-3 variables simultaneously, which was more challenging due to the need to distinguish between multiple options.
+The resulting accuracy levels (Fig 4C, black stars) were significantly higher than chance levels (gray bars).
+We further analyzed the confusion matrix for each variable’s prediction and applied the $\chi^{2}$ test to assess whether the distribution of predictions significantly deviated from the expected distribution under chance.
+We found p-values well below $1\times 10^{-10}$ (see subtitles of Fig 4D, F, G, H, I, J), strongly supporting CREIMBO’s predictive power.
+Furthermore, we assessed the importance of subcircuits’ activations across the four time windows (Fig.29E), with each block of rows representing one subcircuit and the blocks corresponding to different time windows.
+CREIMBO revealed patterns where certain time windows and dynamic coefficients were pivotal for different task variables.
+For instance, $c_{7}$ at $t_{3}$ was critical, while other coefficients, like $c_{9}$ at $t_{3}$, were specific to encoding the lick side.
+This specificity aligns with the fact that this time point likely reflects the final stages of processing during the task, including aspects of learning.
+When we examined $c_{9}$ (Fig.29C), we observed multiple inputs into the secondary motor cortex, related to the execution of the lick towards the trial’s end, highlighting CREIMBO’s ability to pinpoint region-specific interactions.
+Additionally, $c_{8}$ showed increased importance in the initial window $t_{0}$, contributing across multiple variables and capturing inputs into the hippocampus—an area integral to memory processing—underscoring its role in early trial stages.
+These findings further validate CREIMBO’s ability to capture and predict task-related neural dynamics, offering insights into circuit-level brain activity across time.
+Figure 32: Cont. of Figure 11. A: Ground truth latent dynamics ($\{\bm{X}^{d}\}^{D=15}_{d=1}$). B: Ground truth ensemble compositions ($\{\bm{A}^{d}\}^{D=15}_{d=1}$).
+Figure 33: Cont.
+of Figure 11.
+A: Ground truth sub-circuits coefficients ($c$) vary between sessions.
+B: Correlation between ground truth and identified components for A (left), x (middle left), F (middle right), Y (right).
+C: $\ell_{2}$ distance between the components learned by CREIMBO and the ground-truth components for A (left), x (middle left), F (middle right), Y (right).
+Appendix G Inference Assuming Poisson Statistics
+CREIMBO, as outlined in the main text, assumes Gaussian-distributed i.i.d noise.
+This assumption is reflected in the inference procedure, which minimizes the $\ell_{2}$ and Frobenius norms (e.g., equation 1, equation 2).
+Particularly, these norms arise from minimizing the negative log-likelihood of Gaussian-distributed observations noise ($p(\bm{Y}^{d}_{n,t})=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{\bm{Y}^{d}_{n,t}-[A^{d}X^{d}]_{n,t})^{2}}{2\sigma^{2}}\right),$ with some noise standard-deviation $\sigma$.
+Electrophysiology, however, records trains of action potentials (spikes) that are often modeled as a Poisson processes.
+Binned spike counts are thus typically approximated as a Poisson random variable.
+However, when firing rates (FRs) are high enough, the Gaussian assumption is an adequate approximation, and thus the Gaussian assumption remains common in data-driven neuroscience models (e.g., Linderman et al.
+(2016); Aoi & Pillow (2018)).
+However, some species (e.g., bats (Allen et al., 2021)) exhibit extremely low FRs, which may require Poisson likelihood functions.
+Thus we propose an extension to CREIMBO that with a Poisson likelihood function for ensemble and traces inference, that can replace the Gaussian model in low spiking-rates scenarios.
+Our Poisson extension first involves transforming the spike-timing data to binned spike counts for all sessions $d=1\dots D$ ($\{\bm{Y}^{d}\}_{d=1}^{D}$).
+Next, we consider the same goal of finding the ensembles $\bm{A}^{d}$ and their traces $(\bm{X}^{T})^{d}$, under the assumption that the data follows a Poisson distribution.
+To simplify, we focus on a single session $d$, with $\bm{y}=(\bm{Y}^{d})^{T}$ ($T$ for transpose), and note that the extension below applies to all sessions.
+In the Poisson model, for each neuron $n=1\dots N$, the likelihood of a certain spike count at time $t$ given a latent rate $\lambda_{t,n}$ is:
+$p(\bm{y_{t,n}}|\lambda_{t,n})=\frac{\lambda_{t,n}^{y_{t,n}}e^{-\lambda_{t,n}}}{y_{t,n}!}$
+where $y_{t,n}!$ denotes the factorial of $y_{t,n}$.
+Here we model the unknown rate $\bm{\lambda}\in\mathbb{R}^{T\times N}$ as the low-dimensional representation captured via the ensembles $\bm{A}^{d}\in\mathbb{R}^{N\times p}$ and their traces $\widetilde{\bm{x}}:=(\bm{X}^{d})^{T}\in\mathbb{R}^{T\times p}$ ($T$ for transpose), i.e., $\bm{\lambda}=\bm{A}^{d}\bm{X}^{d}$, such that the likelihood becomes
+$p(\bm{y}|\bm{\widetilde{x}},\bm{A})=\prod_{n=1}^{N}\prod_{t=1}^{T}p(y_{t,n}|\lambda_{t,n})$
+$=\prod_{n=1}^{N}\prod_{t=1}^{T}p(y_{t,n}|[\bm{\widetilde{x}}\bm{A}^{T}]_{t,n})$
+$=\prod_{n=1}^{N}\prod_{t=1}^{T}\frac{([\bm{\widetilde{x}}\bm{A}^{T}]_{t,n})^{y_{t,n}}e^{-([\bm{\widetilde{x}}\bm{A}^{T}]_{t,n})}}{y_{t,n}!}.$
+This can be trained by iteratively updating $\bm{\widetilde{x}}$ and $\bm{A}$ to minimize the negative log of the above likelihood, which can be represented mathematically as:
+$\{\widehat{\bm{\widetilde{x}}},\widehat{\bm{A}}\}$ $=\arg\min_{\widetilde{\bm{x}},\bm{A}}\left[-log\left(\prod_{n=1}^{N}\prod_{t=1}^{T}\frac{([\bm{\widetilde{x}}\bm{A}^{T}]_{t,n})^{y_{t,n}}e^{-([\bm{\widetilde{x}}\bm{A}^{T}]_{t,n})}}{y_{t,n}!}\right)\right]$
+$=\arg\min_{\bm{\widetilde{x}},\bm{A}}\sum_{n=1}^{N}\sum_{t=1}^{T}\left[[\bm{\widetilde{x}}\bm{A}^{T}]_{t,n}-y_{t,n}\log([\bm{\widetilde{x}}\bm{A}^{T}]_{t,n})\right]$ (4)
+where the logarithm is taken using the natural exponential base.
+Notably, in the above $\arg\min$, we chose to omit the constant term that emerge from taking the logarithm (i.e., $log(y_{t,n}!)$) as it does not affect the argument minimization.
+In this Poisson case, in contrast to the Gaussian, we can no longer solve for $\bm{x}$ using e.g., least squares or LASSO; instead, we will update $\widetilde{\bm{x}}$ via Gradient Descent.
+The first step is to compute the gradient of the cost function in equation 4 with respect to $\bm{\widetilde{x}}$.
+To simplify the calculation, we first notate the two components obtained in equation 4 by the auxiliary terms
+$g_{1}(\bm{\widetilde{x}},\bm{A})=\sum_{t,n}[\bm{\widetilde{x}}\bm{A}^{T}]_{t,n}$
+and
+$g_{2}(\bm{y},\bm{\widetilde{x}},\bm{A})=\sum_{t,n}-y_{t,n}\cdot\log([\bm{\widetilde{x}}\bm{A}^{T}]_{t,n}).$
+The cost function from equation 4 thus can be written in terms of these functions as:
+$\{\widehat{\bm{\widetilde{x}}},\widehat{\bm{A}}\}=\arg\min_{\bm{\widetilde{x}},\bm{A}}\left[g_{1}(\bm{\widetilde{x}},\bm{A})+g_{2}(\bm{y},\bm{\widetilde{x}},\bm{A})\right].$ (5)
+Hence, for updating $\bm{\widetilde{x}}$ via Gradient Descent, we need to find the gradients of $g_{1}$ and $g_{2}$.
+We first establish two notations to make subsequent steps clearer:
+  • let $[1]_{(m,k)}$ denote a matrix of ones of size $m\times k$.
+  • let $\delta_{(M,P)_{m,p}}$ denote a matrix of shape $M\times P$ whose entries are all zeros except for the entry at index $(m,p)$, which is set to 1.
+Calculate $\frac{\partial g_{1}(\bm{\widetilde{x}},\bm{A})}{\partial\bm{\widetilde{x}}}$:
+To begin, we can rewrite $g_{1}(\bm{\widetilde{x}},\bm{A})$ as
+$g_{1}(\bm{\widetilde{x}},\bm{A})=\sum_{t,n}[\bm{\widetilde{x}}\bm{A}^{T}]_{t,n}=[1]_{(1,T)}\bm{\widetilde{x}}\bm{A}^{T}[1]_{(N,1)}$
+and follow the identity (taken from Petersen et al. (2008) Eq. 20):
+$\frac{\partial[\bm{a}^{T}\bm{M}\bm{b}]}{\partial\bm{M}}=\bm{a}\bm{b}^{T}\hskip 5.0pt\text{ where }\bm{a},\bm{b}\text{ vectors, and }\bm{M}\text{ matrix}.$ (6)
+Building on this, let $[1]_{(1,T)}$ correspond to the $\bm{a}^{T}$ vector from equation 6, $\bm{A}^{T}[1]_{(N,1)}~{}\in~{}\mathbb{R}^{p\times 1}$ correspond to the $\bm{b}$ vector of equation 6, and $\bm{\widetilde{x}}$ correspond to the matrix $\bm{M}$ Then, the gradient of $g_{1}(\bm{\widetilde{x}},\bm{A})$ with respect to $\bm{\widetilde{x}}$ is:
+$\frac{\partial g_{1}(\bm{\widetilde{x}},\bm{A})}{\partial\bm{\widetilde{x}}}=\frac{\partial\left([1]_{(1,T)}\bm{\widetilde{x}}\bm{A}^{T}[1]_{(N,1)}\right)}{\partial\bm{\widetilde{x}}}=[1]_{(T,1)}[1]_{(1,N)}\bm{A}=[1]_{(T,N)}\bm{A}$
+Calculate $\frac{\partial g_{2}(\bm{y},\bm{\widetilde{x}},\bm{A})}{\partial\bm{\widetilde{x}}}$:
+First, we will rewrite the term $[\bm{\widetilde{x}}\bm{A}^{T}]_{t,n}$ as:
+$[\bm{\widetilde{x}}\bm{A}^{T}]_{t,n}=\sum_{j=1}^{p}\bm{\widetilde{x}}_{t,j}(\bm{A}^{T})_{j,n}=\delta_{{(1,T)}_{1,t}}\bm{\widetilde{x}}\bm{A}^{T}\delta_{{(N,1)}_{n,1}}$ (7)
+$\delta_{{(1,T)}_{1,t}}$ is a row vector of length $T$ with a value of 1 at index $t$ and 0 elsewhere, and $\delta_{{(N,1)}_{n,1}}$ is a column vector of length $N$ with a value of 1 at index $n$ and 0 elsewhere.
+We will use again the identity from Petersen et al.
+(2008) presented in equation 6, and now mark $\delta_{{(1,T)}_{1,n}}$ with the equation 6’s vector $\bm{a}^{T}$, $\bm{\widetilde{x}}$ is again $\bm{M}$, and $b:=\bm{A}^{T}\delta_{{(N,1)}_{j,1}}$.
+When applying the identity in equation 6 to equation 7, we thus obtain
+$\frac{\partial[\bm{\widetilde{x}}\bm{A}^{T}]_{t,n}}{\partial\bm{\widetilde{x}}}=\delta_{{(T,1)}_{t,1}}\delta_{{(1,N)}_{1,n}}\bm{A}=\delta_{(T,N)_{t,n}}\bm{A}$
+Given the above, we receive
+$\frac{\partial g_{2}(\bm{y},\bm{\widetilde{x}},\bm{A})}{\partial\bm{\widetilde{x}}}$ $=\frac{\partial\left(\sum_{n=1}^{N}\sum_{t=1}^{T}-y_{t,n}\log\left([\bm{\widetilde{x}}{\bm{A}}^{T}]_{t,n}\right)\right)}{\partial\bm{\widetilde{x}}}$ (8)
+$=-\sum_{n=1}^{N}\sum_{t=1}^{T}\frac{y_{tn}}{[\bm{\widetilde{x}}\bm{A}^{T}]_{t,n}}\delta_{(T,N)_{t,n}}\bm{A}$
+$=-\sum_{t,n}\psi_{t,n}\delta_{(T,N)_{t,n}}\bm{A}$
+$=-\left(\sum_{t,n}\psi_{t,n}\delta_{(T,N)_{t,n}}\right)\bm{A}$
+$=-\bm{\Psi}\bm{A}$
+where $\psi_{t,n}=\frac{y_{t,n}}{[\bm{\widetilde{x}}\bm{A}^{T}]_{t,n}}$, and $\bm{\Psi}\in\mathbb{R}^{T\times N}$ is a matrix whose entry at index $(t,n)$ is $\psi_{t,n}$.
+Going back to the expression in equation 5, the full gradient of the Poisson loss with respect to $\bm{\widetilde{x}}$ is:
+$\frac{\partial[g_{1}(\bm{\widetilde{x}},\bm{A})+g_{2}(\bm{y},\bm{\widetilde{x}},\bm{A})]}{\partial\bm{\widetilde{x}}}=[1]_{(T,N)}\bm{A}-\bm{\Psi}\bm{A}=([1]_{(T,N)}-\bm{\Psi})\bm{A},$
+and the projected gradient step in $\bm{\widetilde{x}}$ at iteration $m$ thus follows:
+$\bm{\widetilde{x}}^{m+1}\leftarrow\Pi\left(\bm{\widetilde{x}}^{m}-\eta^{m}\left([1]_{(T,N)}-\bm{\Psi}\right)\bm{A}\right)$
+where $\Pi$ can project the columns of $\bm{\widetilde{x}}$ onto some constraint set, with the identity transform used if no projection is desired, and $\eta^{m}$ is the gradient descent step size at iteration $m$.
+For inferring $\bm{A}$, at each iteration $m$, we can then use the Spiral-Tap package Harmany et al. (2012) to infer the sparse ensembles.
+The above ensemble model (including $\bm{A}$ and $\bm{x}$) is the only part that directly connects to the Poisson-distributed spike-counts ($\bm{Y}^{d}$), while the latent evolution of $\bm{x}$ via $x_{t}=\bm{F}_{t}\bm{x}_{t-1}$ is independent of the Poisson assumption on the spike-counts $\bm{Y}^{d}$.
+Particularly, as the transition from $\bm{x_{t}}$ to $\bm{x_{t+1}}$ is dictated by the ensembles’ dynamic interactions ($\bm{F_{t}}=\sum_{k=1}^{K}c_{kt}\bm{f}_{k}$), which we continue to assume follows a Gaussian distribution in the latent space, the dynamic inference part of CREIMBO remains unchanged from the main text.